### v1

# An Investigation of FP8 Across Accelerators for LLM Inference

Abstract
The introduction of 8-bit floating-point (FP8) computation units in modern AI accelerators has generated significant interest in FP8-based large language model (LLM) inference. Unlike 16-bit floating-point formats, FP8 in deep learning requires a shared scaling factor. Additionally, while E4M3 and E5M2 are well-defined at the individual value level, their scaling and accumulation methods remain unspecified and vary across hardware and software implementations. As a result, FP8 behaves more like a quantization format than a standard numeric representation. In this work, we provide the first comprehensive analysis of FP8 computation and acceleration on two AI accelerators: the NVIDIA H100 and Intel Gaudi 2. Our findings highlight that the Gaudi 2, by leveraging FP8, achieves higher throughput-to-power efficiency during LLM inference, offering valuable insights into the practical implications of FP8 adoption for datacenter-scale LLM serving.
FP8, LLM, Quantization, GEMM, GEMV
1 Introduction
The FP8 format has emerged as a promising datatype due to its lower computational overhead, reduced power consumption, and improved memory efficiency compared to commonly used formats such as FP32 and BF16 (Kalamkar et al., 2019), particularly at large batch sizes. With the introduction of the Intel Gaudi HPU and NVIDIA Hopper GPUs, FP8 is gaining traction as a key datatype for next-generation AI workloads (DeepSeek-AI, 2024).
While extreme low-bit quantization has been widely studied, its throughput remains inferior to conventional 16-bit floating-point formats on commercial hardware (Lin et al., 2024b). In contrast, FP8 has rapidly become the preferred format due to its compatibility with existing accelerators while preserving model accuracy. This makes it particularly well-suited for datacenter-scale LLM inference, where both high throughput and minimal accuracy degradation are critical. In some accelerators, FP8 can be twice as fast as BF16. Given the immense computational costs of training and deploying large language models (LLMs), FP8 has garnered significant attention for its potential efficiency gains.
Despite its advantages, FP8 utilization is constrained by the widening gap between computing speed and memory bandwidth. A key challenge is that, while LLM sizes continue to scale, their hidden dimensions grow at a much slower rate (e.g., Llama 8B has a hidden size of 4096, but Llama v3.3 70B only increases hidden size to 8192). Tensor parallelism (TP) further partitions matrices into even smaller fragments, making small matrix efficiency a critical performance factor.
Especially during the decode phase, Generalized Matrix-Vector (GEMV) and thin General Matrix-Matrix Multiply (GEMM) operations dominate, reducing Model FLOP Utilization (MFU) (Chowdhery et al., 2023; Narayanan et al., 2021). With the growing popularity of “reasoning” models (DeepSeek-AI, 2025), the proportion of decoding in LLM inference workloads can be expected to continue growing.
To fully address the challenges of LLM inference, we set out to quantify the various constraints involved, including those that arise with the inclusion of FP8 quantization. To the best of our knowledge, we are the first to provide quantitative measurements associated with various FP8 quantization schemes.
There is a notable lack of research providing comprehensive FP8 experimental results, particularly in studies analyzing hardware characteristics in LLM inference configurations. To address this gap, our work offers a detailed evaluation of FP8 performance across different hardware platforms, providing valuable insights for model developers on the trade-offs between model architecture and hardware design. To summarize our contributions:
Figure 1: TFLOPS (top row) and TFLOPS/Watt (bottom row) comparison between the Gaudi 2 and H100 using FP8 in the decode phase with batch size set to 64. We measure inference throughput using Optimum Habana v1.15.0 for the Gaudi 2 and TensorRT-LLM v0.16.0 for the H100. For short sequences, the Gaudi 2 has an advantage both in absolute throughput and in TFLOPS/Watt. At longer sequences, the H100 achieves higher TFLOPS due to its superior memory bandwidth, but is less energy efficient due to higher power consumption.
  1. 1.
    We analyze scaling factor configurations for FP8 formats and quantify key LLM inference constraints in datacenter environments, providing a unified framework to compare throughputs across different inference stages, model sizes, and sequence lengths.
  2. 2.
    We analyze the computational characteristics of LLM inference on the Gaudi 2 and H100 across both the prefill and decode phases. In the prefill phase, where compute performance is the key factor, H100’s throughput is 50% higher than Gaudi 2’s. However, since the decode phase dominates overall inference time, accelerator utilization becomes more critical, and Gaudi 2 achieves 50% higher throughput in this phase. This highlights the differing performance trade-offs between the two accelerators across LLM inference workloads.
2 Background
With the explosive growth of LLM sizes in recent years, numerous techniques specific to generative LLMs have emerged to reduce memory requirements and improve output speeds. In this work, we focus on the large-batch serving environment often found in data centers, where multiple requests for LLM inference run simultaneously under power constraints at both the rack and data center levels.
Various methods have been proposed to improve LLM serving efficiency, with quantization being one of the most widely studied. However, a key challenge in LLM quantization is the presence of activation outliers. Transformer models often contain large outlier values in their activations, which, if not properly addressed during quantization, can lead to substantial quality degradation (Sun et al., 2024a). To mitigate this issue, several quantization techniques have been developed, aiming to improve robustness and accuracy while maintaining computational efficiency.
2.1 Weight-only quantization
Due to the quantization difficulties posed by activation outliers, many studies have proposed several weight-only quantization (WoQ) methods(Lin et al., 2024a; Frantar et al., 2023; Lee et al., 2023), while retaining activations in BF16 or FP16 format. However, these approaches often suffer from low throughput at larger batch sizes. To enable computations, quantized weights must be converted into the same data type as activations (e.g. BF16), requiring online dequantization. This is handled by CUDA cores on NVIDIA GPUs and TPC cores on Gaudi HPUs, which operate an order of magnitude slower than their respective matrix multiplication units. Therefore, dequantization frequently becomes a bottleneck even at moderate batch sizes. Also, because GEMM computations are conducted at higher precisions, WoQ methods are fundamentally limited to having the same throughput as unquantized GEMMs at best.
Even when high throughput is achieved, kernel implementations often need to be re-implemented for different devices, including those from the same vendor. For instance, the Marlin kernel (Frantar et al., 2024), optimized for int4-weight and float16-activations on the A100, is suboptimal for the H100. As a result, the Machete kernel (Wilkinson, 2024) was developed specifically to leverage the new features of the Hopper architecture, employing different design trade-offs. These challenges arise due to variations in cache sizes, hardware features, and compute-to-memory bandwidth ratios, all of which necessitate different optimizations to achieve high performance. Consequently, adapting WoQ methods across different accelerators imposes significant implementation overhead and increases development costs.
2.2 INT8 quantization
The INT8 quantization format described in Xiao et al. (2024) is closely related to our work, requiring quantization of both weights and activations to leverage INT8 tensor cores to accelerate computations. While INT8 quantization has been widely adopted in fields such as computer vision, its limited dynamic range presents challenges in achieving full accuracy for LLM inference. FP8 thus has an advantage over INT8 as integer representations are inherently less robust to outliers compared to floating-point numbers (Shen et al., 2024; Kurtic et al., 2024).
2.3 FP8 quantization
The FP8 format can be categorized into two variants based on the allocation of bits to the exponent and mantissa: the E4M3 format and the E5M2 format. E4M3 offers higher precision compared to the E5M2 format, which instead has a higher dynamic range. To address the smaller dynamic range of FP8 compared to BF16, scaling factors are employed to adjust the inputs before casting to FP8. This ensures that the values fit within the FP8 range, maximizing the utilization of its limited dynamic range. Per-tensor scaling, which computes a single scaling factor for each matrix. For more granular control, per-token scaling can be applied to activations, while per-channel scaling is used for weights, enabling finer adjustments. Group-wise scaling factors (DeepSeek-AI, 2024) have also been proposed.
2.4 Scaling factor computation
A critical consideration for activation quantization is the overhead of converting activations to lower precision. While non-linear operations require 16-bit or even 32-bit precision, activations must be quantized online immediately before GEMM computations utilize low-precision compute units. This introduces non-trivial overhead, limiting the throughput gains from reduced precision. Activation quantization also requires estimating input statistics to determine appropriate scaling factors. A direct approach, measuring statistics for each activation before quantization, would necessitate two memory accesses: one for collecting statistics and another for quantization, incurring high latency. To mitigate this, many libraries use a calibration set to precompute scaling factors based on a representative sample of activations.
A key limitation of static scaling is that it typically enforces per-tensor scaling of activations, as token-wise variations are unknown in advance. For instance, altering the order of sequences during batched decoding would shift token positions. Moreover, using position-dependent scaling factors would require fixed input sizes. While per-channel scaling (Xiao et al., 2024) is possible, the same scaling factors are still applied to all tokens, potentially reducing accuracy.
In contrast, dynamic scaling is more efficient when applied separately to each token as measurement and quantization can be combined in a single memory access. It also enables group-wise quantization, as implemented in DeepSeek-AI (2024) with a group size of 128, improving output quality through finer-grained scaling factors.
A potential optimization for both scaling methods is to merge quantization with other memory-bound operations. In Pre-LN Transformers (Xiong et al., 2020), activation quantization can be fused with RMSNorm. Similarly, element-wise multiplication in SwiGLU (Shazeer, 2020) can also be fused. Frameworks such as TensorRT-LLM (NVIDIA, 2023) v0.16.0 have already implemented these optimizations for FP8 activation quantization. They have yet to be implemented for Gaudi HPUs as of the time of writing.
2.5 LLM evaluation
LLM evaluation has grown in importance as new LLMs compete for prominence. Traditionally, WikiText-2 perplexity (Merity et al., 2017) has been widely used due to its ease of measurement and relationship with the training loss. More recently, Massive Multitask Language Understanding (MMLU) (Hendrycks et al., 2021) has emerged as the de facto standard to assess LLM quality, evaluating knowledge in a wide range of domains. However, multiple-choice benchmarks may fail to capture the full impact of quantization. For instance, the Llama v3 405B model (Llama Team, 2024) applies max-absolute (maxabs) clipping when calculating scaling factors despite there being no effect in standard evaluation metrics because the model may generate corrupted outputs for high-perplexity tokens, such as dates, without maxabs clipping.
Moreover, evaluation should extend beyond pre-trained base models to instruction-tuned models that have undergone alignment. If quantization is applied after pre-training but before alignment, the alignment process must be quantization-aware, requiring each quantization algorithm to be integrated with each alignment method while maintaining alignment quality, significantly increasing complexity. For this reason, we conduct all evaluations on instruction-t uned models, where quantization effects are more representative of real-world deployment scenarios.
3 Matrix multiplication comparison
3.1 Differences in hardware capabilities
We compare implementations of FP8 GEMM between NVIDIA GPUs and Intel Gaudi HPUs, identifying key differences despite both adhering to the FP8 specification (Micikevicius et al., 2022).
(Accumulation precision) Hopper GPUs use a 14-bit accumulator for FP8 GEMMs (DeepSeek-AI, 2024), requiring casting to CUDA cores for higher precision. Software optimizations, such as applying FP32 accumulation to only one in four WGMMA instructions, reduce error but increase kernel complexity and remain less precise than full FP32 accumulation. In contrast, Gaudi HPUs always accumulate in FP32, ensuring higher numerical precision.
(E4M3 range) The Gaudi 2 follows the IEEE specification for special values, unlike NVIDIA GPUs, which use a single special value representation (Noune et al., 2022). This results in seven fewer magnitude representations and a maximum value of 240 for E4M3 in the Gaudi 2 compared to 448 on NVIDIA GPUs. This has been addressed in the Gaudi 3, but we were unable to test these in our experiments.
(Power-of-2 scaling) Gaudi HPUs allow modification of floating-point exponent biases to accelerate scaling factor application. The Gaudi 2 supports fixed hardware-accelerated scaling factors of $2^{-8},2^{-4},2^{0},2^{4}$ for E4M3 while the Gaudi 3 extends this to arbitrary powers of 2 between $2^{-32}$ and $2^{31}$. However, this feature is limited to GEMMs with per-tensor scaling factors.
(Stochastic rounding) During FP8 quantization, stochastic rounding can be applied when converting 16/32-bit floating-point values to FP8, similar to the technique proposed for FP32-to-BF16 conversion in (Fan et al., 2024). This method is distinct from stochastic rounding applied at the inner-product (El Arar et al., 2023).
(Sparsity) NVIDIA GPUs support semi-structured sparsity acceleration, potentially doubling tensor core peak throughput. However, despite attempts to leverage sparsity in LLMs (Frantar & Alistarh, 2023; Sun et al., 2024b), dense GEMMs remain dominant due to accuracy loss and limited speedups. Gaudi HPUs do not support sparsity acceleration.
3.2 GEMM throughput measurements
Table 1: Throughput and power measurements for square FP8 GEMMs with row-wise scaling. The ratio of measured TFLOPS and power draw relative to their peak values are included in parentheses. We include the TFLOPS/Watt ratio on the right. H100 has 1989.9 peak FP8 GEMM TFLOPS and 700W TDP. Gaudi 2 has 865 peak FP8 GEMM TFLOPS and 600W TDP.
Device (M,K,N) TFLOPS Power (W) TF/W
Gaudi 2 1K 367.9 (42.5%) 375 (63%) 1.0
2K 586.2 (67.8%) 460 (77%) 1.3
4K 817.1 (94.5%) 460 (77%) 1.8
8K 741.8 (85.8%) 490 (82%) 1.5
H100 1K 218.3 (11.0%) 350 (50%) 0.6
2K 879.7 (44.2%) 690 (99%) 1.3
4K 1167.6 (58.7%) 690 (99%) 1.7
8K 1084.7 (54.5%) 690 (99%) 1.6
For a GEMM between matrices of size $(M\times K)\times(K\times N)$, the total number of floating-point operations (FLOPs) performed is $2MKN$. This is derived from the $M\times N$ dot products of length $K$, where each element undergoes one multiplication and one addition. Following the convention where FLOPS denotes FLOPs per second, we compute throughput in FLOPS using the theoretical FLOPs and the measured latency. We can thus calculate the MFU by dividing observed throughput by peak throughput.
As LLM computations are dominated by matrix multiplications, GEMM throughput serves as an upper bound for model end-to-end performance. While accelerator specifications list peak throughput values, actual throughput rarely reaches this limit, particularly for small matrices.
Table 1 presents GEMM throughput and power consumption measurements for row-wise scaled FP8 GEMM. We conducted measurements on NVIDIA H100 GPUs using the NGC PyTorch 24.12 image and on Intel Gaudi 2 HPUs using the Synapse AI v1.19.0 image. Power consumption was recorded via the ”nvidia-smi” and ”hl-smi” utilities, respectively. Input casting overheads were excluded from the measurements. Additional results for GEMM throughputs under different conditions are provided in Tables 8 and 9. Our results show that the Gaudi 2 achieves higher utilization than the H100, particularly for small matrices. The performance drop for smaller matrices is steeper on the H100, with the Gaudi 2 providing a higher TFLOPS at matrix sizes of 1K. Also, the Gaudi 2 exhibits lower power consumption than its stated 600W TDP, whereas the H100 approaches its peak 700W TDP even at moderate utilizations.
Unlike NVIDIA GPUs, which incorporate multiple small matrix multiplication accelerators, Gaudi HPUs feature a few large matrix accelerators. As a result, Gaudi HPUs requires fewer input elements per cycle to fully utilize compute resources, thereby reducing first-level memory bandwidth requirements and improving efficiency (Intel, 2022, 2024). Additionally, Gaudi HPUs leverage a graph compiler to dynamically reconfigure the MME size based on the target GEMM dimensions, optimizing resource utilization.
This superior efficiency of the Gaudi 2 for small matrices is particularly relevant given that LLM hidden dimensions typically range between 1K (e.g., Llama 1B) and 8K (e.g., Llama 70B), with tensor parallelism further reducing matrix sizes. Moreover, the lower power draw of the Gaudi 2 relative to its TDP suggests that naïve TDP comparisons can be misleading, emphasizing the need for empirical evaluation.
4 Comparing end-to-end results
We conduct evaluations on different FP8 features and come to the following conclusions. First, E4M3 consistently outperforms E5M2 in quantization accuracy across nearly all tested configurations. Second, stochastic rounding during quantization has minimal impact on model accuracy and, in some cases, may even be detrimental. Third, dynamic scaling achieves accuracy comparable to BF16 models while eliminating the need for a calibration set.
For evaluation, we use instruction-tuned versions of Llama v3.2 1B, v3.2 3B, v3.1 8B, and v3.3 70B. Unless stated otherwise, models employ dynamic row-wise scaling for all linear layers, while LM head remains in BF16.
Table 2: The “Cited” columns include results from Kim (2025) using static “FP8 (S)” scaling from the Intel Neural Compressor (INC) library using the UltraChat 200K (Ding et al., 2023) dataset for calibration. The “Measured” column includes results for dynamic “FP8 (D)” scaling. Reference BF16 accuracies are included for both columns to control for evaluation condition differences.
Llama v3.1 8B Inst. Cited Measured
BF16 FP8 (S) BF16 FP8 (D)
MMLU CoT 5-shot 68.4% 66.3% 68.8% 68.3%
GSM8K CoT 5-shot 75.7% 70.4% 83.1% 84.5%
Winogrande 78.1% 77.6% 73.8% 73.8%
TruthfulQA mc1 37.1% 34.9% 39.9% 39.4%
TruthfulQA mc2 54.0% 52.2% 55.1% 54.3%
4.1 Dynamic vs static scaling
Section 2 describes the trade-offs between static and dynamic scaling for activation quantization. Static scaling offers higher throughput for large matrices as scaling factors do not need to be computed for each input. It also enables per-tensor quantization, leading to improved GEMM utilization. In contrast, dynamic scaling enhances output quality by assigning separate scaling factors to each token.
Table 2 presents a comparative analysis of model output quality on static vs. dynamic FP8 scaling. The results indicate that dynamic quantization obtains results similar to the original BF16 models, whereas static quantization leads to noticeable accuracy degradation. While various techniques, such as those proposed by Xiao et al. (2024); Liu et al. (2024); Ashkboos et al. (2024), can potentially mitigate this degradation, they still pose a risk as the scaling factors obtained during calibration may be suboptimal for unseen inputs. Therefore, dynamic quantization is the preferred approach, provided that the associated throughput reduction remains within acceptable limits.
4.2 E4M3 vs. E5M2
Table 3: Comparison between different FP8 data types and rounding modes for MMLU CoT 5-shot performance on instruction-tuned Llama models. Stochastic rounding provides little or no benefit to accuracy while E5M2 is detrimental, especially for smaller models. Experiments were conducted on a Gaudi 2 HPU.
Model Data Type Rounding MMLU
Llama v3.2 BF16 - 46.3%
1B Instruct E4M3 SR 45.7%
E4M3 RTN 45.5%
E5M2 RTN 44.5%
Llama v3.2 BF16 - 61.8%
3B Instruct E4M3 SR 61.7%
E4M3 RTN 61.6%
E5M2 RTN 60.7%
Llama v3.1 BF16 - 68.8%
8B Instruct E4M3 SR 68.3%
E4M3 RTN 68.3%
E5M2 RTN 67.5%
Llama v3.3 BF16 - 82.0%
70B Instruct E4M3 SR 82.0%
E4M3 RTN 82.0%
E5M2 RTN 82.2%
Following on the findings of Shen et al. (2024), which demonstrated that the E4M3 format achieves superior accuracy on language tasks compared to E5M2, we extend this analysis to instruction-tuned models in Table 3. To determine the optimal format for inference, we conducted a comparative evaluation of E4M3 and E5M2, measuring MMLU accuracy using LM Evaluation Harness (v0.4.7, (Gao et al., 2024)). Our experimental results consistently indicate that E4M3 outperforms E5M2 across all evaluated scenarios. Furthermore, as shown in Table 8, the GEMM throughputs using E4M3 and E5M2 are comparable, making E4M3 the preferred choice for inference even considering the smaller representational capacity of E4M3 on Gaudi 2.
Figure 2: Process and utilization characterization of the two phases of generative LLM inference.
4.3 Stochastic rounding
Hardware-accelerated stochastic rounding during FP8 quantization is a distinctive feature of Gaudi HPUs, similar to the approach proposed by Fan et al. (2024), and is unavailable in NVIDIA GPUs. Equation 1 formalizes the concept, where a higher-precision value $x$ is rounded up to $x_{up}$ or down to $x_{down}$ stochastically based on the distance from $x$.
$x_{quantized}=\begin{cases}x_{up}\quad(p=\frac{x-x_{down}}{x_{up}-x_{down}})\\ x_{down}\quad(p=\frac{x_{up}-x}{x_{up}-x_{down}})\end{cases}$ (1)
Stochastic rounding is expected to preserve more of the original numerical information post-quantization, potentially leading to improved model accuracy. However, empirical results in Table 3 indicate that stochastic rounding during quantization does not significantly enhance model accuracy. Furthermore, as shown in Table 6, it may even lead to accuracy degradation in certain cases. These findings suggest that while stochastic rounding during quantization theoretically retains more information, its practical benefits for FP8 inference in LLMs remains inconclusive.
5 LLM inference comparison
5.1 LLM inference phases
Generative LLM inference consists of a compute-bound prefill phase and a memory-bound decode phase (Patel et al., 2024; Zhong et al., 2024). We include a diagram highlighting the differences between the two phases in Figure 2.
In the prefill phase, the LLM generates the first token based on the provided input prompt. The key-value pairs are stored in the KV cache for reuse during the generation of subsequent tokens. The prefill phase processes the entire input in parallel, utilizing matrix-matrix operations that are typically compute-bound. This parallelism allows for effective hardware utilization, enabling efficient input processing.
After the first token is generated, the process transitions to the decode phase, where output tokens are generated sequentially in an autoregressive manner based on the previous tokens. The decode phase involves GEMV operations, inherently limiting hardware utilization compared to the compute-bound GEMMs in the prefill phase.
Modern inference frameworks such as vLLM (Kwon et al., 2023) and TensorRT-LLM (NVIDIA, 2023) partially remedy this issue by batching multiple decoding requests, improving the computational intensity of the linear layers. However, the batch size is limited by the memory capacity as each sequence in a batch requires its own KV cache. Furthermore, the computational intensity of the attention operation during decoding is unaffected by batch size, instead requiring a GEMV for each sequence. Grouped query attention (GQA) (Ainslie et al., 2023) converts the operation into a thin GEMM, but it remains memory-bound. This slowdown becomes more evident at longer sequence lengths, where the attention FLOPs continue to increase in proportion to sequence length while the FLOPs required for linear layers remain constant.
5.2 Calculating inference FLOPs
While previous works (Patel et al., 2024) have evaluated inference performance using time to first token (TTFT) and time per output token (TPOT), these metrics do not facilitate comparisons across different inference stages, model sizes, or sequence lengths. To ensure a consistent comparison, we directly compute model FLOPs. Following the approach of Ansel et al. (2024); Narayanan et al. (2021), we calculate only the FLOPs associated with matrix multiplication, as these dominate LLM computations. However, unlike previous works, we exclude FLOPs related to autoregressive attention masking, which can be skipped (Dao, 2024). This approach aligns more closely with how the KV cache is leveraged when generating new tokens.
Using this method, the FLOPs required for a forward pass of a Llama model with $l$ transformer blocks, hidden size $h$, intermediate size $ah$, head size $d$, head count $H=h/d$, GQA group size $g$, vocabulary size $v$, and input sequence of length $s$ can be calculated as follows:
$f_{llama}(s)=2sh^{2}l(3a+2+\frac{2}{g})+2s^{2}hl+2vsh$ (2)
By denoting $A=(3a+2+\frac{2}{g})$ as a constant for each model, the equation can be simplified to the following:
$f_{llama}(s)=2s(Ah^{2}l+vh)+2s^{2}hl$ (3)
When the model generates $t$ tokens in a single decoding iteration where $t\ll s$, we can calculate Equation 2 using $s^{\prime}=s+t$ and obtain the approximation:
$f_{llama}(s+t)-f_{llama}(s)$ (4)
$\approx 2t(Ah^{2}l+vh)+4sthl$
From Equation 4, we observe that linear layer computations, including those in the LM head, remain independent of the previous sequence length $s$. However, attention computations scale proportionally with $s$. In the autoregressive case where $t=1$ and considering a batch of sequences with batch size $b$ and sequence lengths $s_{1},...,s_{b}$, the FLOPs per decoding step can be approximated as follows:
$2b(Ah^{2}l+vh)+4hl\sum_{i=1}^{b}s_{i}$ (5)
One challenge in interpreting these results is that only the $2bAh^{2}l$ term, representing linear layers except the LM head, is computed in FP8, whereas the $2bvh$ term for the LM head and the $4hl\sum_{i=1}^{b}s_{i}$ term for attention are computed in BF16. Another is that online KV cache dequantization would add non-trivial overhead to Equation 5. However, such overheads do not constitute additional model FLOPs as per the definition of MFU in Chowdhery et al. (2023).
A more significant limitation is that each FLOP cannot be executed at full efficiency due to hardware utilization constraints. For example, the Gaudi 2 has a peak HBM bandwidth of 2.4 TB/s and a peak FP8 GEMM throughput of 865 TFLOPS, requiring a FLOP/byte ratio (computational intensity, CI) of at least 360 for optimal FP8 execution. However, in the decoding phase, a thin GEMM of $(b\times h)\times(h\times ah)$, where $b\ll h$ and $a\geq 1$, results in a CI of approximately $2b$ for FP8 and $b$ for BF16, far below the required intensity for peak throughput on realistic batch sizes $b$. Additionally, matrix multiplication units operate with fixed block sizes, meaning that full utilization is only achieved when input dimensions align with hardware-friendly shapes such as multiples of 128 (Lee et al., 2024).
Furthermore, KV cache computations present a unique bottleneck. Increasing batch size does not improve the CI since each sequence in a batch maintains a separate KV cache. For a BF16 KV cache using GQA with $g$ groups, the CI is thus limited to $g$ FLOPs/byte. Consequently, even with a perfectly optimized kernel, the theoretical throughput of applying the query to the KV cache is capped by the memory bandwidth multiplied by the CI. For Llama v3 models with $g=8$ on a Gaudi 2 with a peak 2.4 TB/s HBM, this theoretical ceiling is 19.2 TFLOPS, a small fraction of peak GEMM throughput. As attention computations are both inherently memory-bound and scale linearly with sequence length, decoding at longer sequence lengths ultimately converges to the attention throughput, as demonstrated in Figure 1.
By distinguishing these phases and understanding their computational characteristics, optimization strategies can be tailored to enhance the efficiency of LLM inference, especially during the resource-intensive decode phase.
5.3 Prefill phase
Figure 3: Roofline diagram during the prefill phase for batch size 1. Throughput improves with longer sequence length until it begins to decline as the proportion of attention computations, which are slower than GEMMs, take up a greater share of the computation. We use static FP8 scaling for both the H100 and the Gaudi 2 as optimized dynamic FP8 scaling is unavailable for the Gaudi.
We first conducted an analysis of the operational characteristics of the two accelerators in the prefill phase. Figure 3 compares the prefill TFLOPS achieved for three models across varying sequence lengths, using a roofline diagram for the H100 and Gaudi 2 accelerators. The results demonstrate that the H100 achieves significantly higher throughput during the prefill phase than the Gaudi 2. For Llama 8B, the H100 achieves double the throughput of the Gaudi 2.
During the prefill phase, attention layers process long input sequences, enabling large matrix computations to be executed in a single step. This maximizes compute engine utilization, leading to high efficiency. As a result, performance in this phase is primarily dictated by GEMM throughput, rather than utilization or memory bandwidth constraints.
As model sizes increase, matrix dimensions grow proportionally with hidden dimensions, making computations more compute-bound. This leads to a clear trend of increasing prefill throughput with larger models, underscoring the importance of compute performance in the prefill phase. A similar pattern is observed for longer sequence lengths, where utilization improves until reaching saturation.
Figure 4: Decode throughput comparison between BF16 and FP8 on Llama v3.1 8B Instruct using a batch size of 64 using static FP8 scaling on the Gaudi 2 (left); static and dynamic FP8 scaling on the H100 (right). Throughput differences between FP8 and BF16 in the Gaudi 2 are 50% or greater, whereas they are under 25% for the H100. Dynamic scaling has higher throughput despite its greater overhead because row-wise GEMM is faster than per-tensor GEMM for small matrices on the H100 (see Table 9).
5.4 Decode phase
In Figure 1, we compare decode throughputs between the Gaudi 2 and H100. Surprisingly, the Gaudi 2 not only demonstrates superior power efficiency across all sequence lengths but also achieves higher absolute throughput for short sequences below 4K tokens. At longer sequence lengths, however, memory bandwidth becomes the primary bottleneck. Our measurements show HBM bandwidths of 2.0 TB/s for the Gaudi 2 and 2.6 TB/s for the H100, giving the latter an advantage as the KV cache size increases.
We also compare decode throughputs for BF16, dynamic FP8 scaling, and static FP8 scaling on the H100 in Figure 4 and find that the throughput gain from FP8 is below 25%. In contrast, the gain for the Gaudi 2 approaches 50%.
Figure 5: Thin GEMM MFU comparison between the Gaudi 2 and H100 for both BF16 and FP8. The Gaudi 2 maintains a similar MFU for BF16 and FP8 of similar shapes while there is a noticeable drop for the H100. The MFU differences at the same shapes are enough to provide superior TFLOPS for the Gaudi 2 over the H100 as shown in Table 4.
To investigate the lack of performance gains in the H100, we analyze its throughput on thin matrices, identifying it as the primary bottleneck. Table 4 shows that while FP8 GEMM throughput on the Gaudi 2 is nearly twice that of BF16 GEMM, there is minimal improvement on the H100. As shown in Equation 5, the decoding stage at short sequence lengths ($s<h$) is dominated by linear operations, making thin GEMM performance critical.
Table 4: Thin GEMM throughputs in TFLOPS for the Gaudi 2 and H100 measuring $(M\times K)\times(K\times N)$ GEMMs, where $M$ corresponds to typical batch sizes encountered during inference and $K,N$ represent hidden dimension sizes. Row-wise scaling is used for FP8 GEMMs. Throughput scales linearly with $M$ on both devices, but the Gaudi 2 consistently outperforms the H100, even for BF16. These results highlight the superior efficiency of the Gaudi 2 for thin GEMMs, crucial for LLM decoding.
GEMM TFLOPS Gaudi 2 H100
Shape: (M,K,N) BF16 FP8 BF16 FP8
( 8, 1024, 1024) 3.3 3.8 1.7 1.7
(16, 1024, 1024) 6.5 11.4 3.4 3.9
(32, 1024, 1024) 12.8 23.8 6.5 7.0
(64, 1024, 1024) 26.7 54.0 12.6 14.9
( 8, 2048, 2048) 12.4 26.1 6.7 7.5
(16, 2048, 2048) 20.6 48.6 12.9 15.0
(32, 2048, 2048) 48.0 87.6 27.1 28.2
(64, 2048, 2048) 91.3 163.2 52.3 60.5
( 8, 4096, 4096) 18.8 35.4 14.4 16.8
(16, 4096, 4096) 37.4 67.9 28.6 33.5
(32, 4096, 4096) 73.6 132.0 68.3 68.1
(64, 4096, 4096) 144.5 253.4 133.3 133.9
6 Conclusion
In this study, we provide a comprehensive analysis of FP8, an emerging numerical format that remains underexplored. We examine various FP8 scaling factor configurations and analyze LLM inference workloads based on the constraints of real-world data center environments. We make a clear exposition of the key constraints in LLM inference, also providing a framework to make precise calculations for inference throughput. Our analysis of the two phases of LLM inference highlights the strong dependence of LLM inference on hardware characteristics.
However, beyond hardware development, software optimizations that fully leverage accelerator capabilities are equally essential to maximizing performance. One key limitation of the Gaudi is its limited software support and integration. For instance, many kernel fusion techniques for dynamic FP8 scaling are currently unavailable on the Gaudi. Additionally, the decode throughputs reported in our study relied on static memory allocation as the paged attention implementations remained uncompetitive despite recent optimizations (Lee et al., 2024). Even for static memory allocation, the observed attention throughputs are lower than what would be expected from the hardware specifications. Furthermore, while not a primary focus of this work, the lack of KV cache quantization support present challenges for deployment.
Despite these concerns, the growing adoption of inference-time scaling via chain-of-thought (CoT) (Wei et al., 2022) prompting suggests that LLM workloads will increasingly be dominated by decoding. Emerging techniques that decouple the prefill and decode phases (Patel et al., 2024; Zhong et al., 2024) enable hardware to be allocated to the tasks for which they are best suited. Given our findings that the Gaudi 2 demonstrates superior decoding efficiency for FP8, this has potential implications for cloud service providers seeking cost-effective inference solutions.
Appendix A Appendix
Table 5: MMLU result for per-row, per-tensor options
  model Rowwise Scaling accuracy
Activation Weight
Llama-3.2-1B 35.12%
34.69%
34.36%
34.65%
Llama-3.2-3B 53.90%
53.92%
53.83%
53.75%
Llama-3.1-8B 62.95%
62.92%
62.80%
62.90%
Llama-3.1-70B 75.25%
75.23%
75.32%
75.39%
Table 6: MMLU 5-shot CoT for stochastic rounding using E4M3.
  Model Stochastic Rounding MMLU accuracy
Activation Weight
Llama-3.2-1B 33.94%
34.31%
35.20%
34.65%
Llama-3.2-3B 54.09%
53.95%
54.41%
53.75%
Llama-3.1-8B 62.24%
62.43%
62.55%
62.90%
Llama-3.1-70B 75.05%
75.18%
75.30%
75.39%
Table 7: Throughput and Power Measurement of Llama Instruct Models
  Model Name Mode Batch Size Seq Len Power(W) Prefill TFLOPS/HPU
Llama-3.2-1B-Instruct FP8-Static 4 512 509 412
Llama-3.2-1B-Instruct FP8-Static 4 1024 509 411
Llama-3.2-3B-Instruct FP8-Static 4 512 498 518
Llama-3.2-3B-Instruct FP8-Static 4 1024 511 537
Llama-3.1-8B-Instruct FP8-Static 4 512 471 588
Llama-3.1-8B-Instruct FP8-Static 4 1024 436 595
Table 8: Gaudi 2 throughput in TFLOPS for scaled FP8 GEMM for square matrices of the given size, excluding quantization overhead. Measured throughput relative to the peak FP8 throughput (865 TFLOPS) is included in parentheses. Hardware acceleration is only available for per-tensor scaling.
Type Size Per-Row Per-Tensor HW Accel.
E4M3 1K 494 (57.1%) 494 (57.1%) 494 (57.1%)
2K 506 (58.5%) 641 (74.1%) 641 (74.2%)
4K 735 (84.9%) 796 (92.1%) 801 (92.6%)
8K 742 (85.7%) 822 (95.0%) 852 (98.4%)
E5M2 1K 306 (35.4%) 494 (57.1%) 493 (57.0%)
2K 506 (58.5%) 642 (74.2%) 642 (74.2%)
4K 735 (84.9%) 802 (92.7%) 802 (92.7%)
8K 726 (83.9%) 825 (95.4%) 825 (95.4%)
Table 9: H100 throughput in TFLOPS for scaled FP8 GEMM for square matrices of the given size, excluding quantization overhead. Measured throughput as a ratio of peak FP8 throughput (1989.9 TFLOPS) is included in parentheses. We only include results for E4M3 because we were unable to find libraries exposing low-level E5M2 GEMM APIs on NVIDIA GPUs.
Accum. Size Per-Row Per-Tensor
FP32 1K 217 (11.0%) 186   (9.4%)
2K 299 (15.1%) 840 (42.4%)
4K 362 (18.3%) 1099 (55.5%)
8K 396 (20.0%) 1300 (65.7%)
Fast 1K 237 (12.0%) 147   (7.4%)
2K 810 (40.9%) 896 (45.3%)
4K 1136 (57.4%) 1205 (60.9%)
8K 1123 (56.8%) 1388 (70.1%)

### v2

@@
 With the growing popularity of “reasoning” models (DeepSeek-AI, 2025), the proportion of decoding in LLM inference workloads can be expected to continue growing.
-To fully address the challenges of LLM inference, we set out to quantify the various constraints involved, including those that arise with the inclusion of FP8 quantization.
+To comprehensively address the challenges of LLM inference, we aim to quantify the various constraints that impact performance, including those introduced by FP8 quantization.
 To the best of our knowledge, we are the first to provide quantitative measurements associated with various FP8 quantization schemes.
@@
 3.2 GEMM throughput measurements
-Table 1: Throughput and power measurements for square FP8 GEMMs with row-wise scaling.
-The ratio of measured TFLOPS and power draw relative to their peak values are included in parentheses.
-We include the TFLOPS/Watt ratio on the right.
-H100 has 1989.9 peak FP8 GEMM TFLOPS and 700W TDP.
-Gaudi 2 has 865 peak FP8 GEMM TFLOPS and 600W TDP.
-Device (M,K,N) TFLOPS Power (W) TF/W
-Gaudi 2 1K 367.9 (42.5%) 375 (63%) 1.0
-2K 586.2 (67.8%) 460 (77%) 1.3
-4K 817.1 (94.5%) 460 (77%) 1.8
-8K 741.8 (85.8%) 490 (82%) 1.5
-H100 1K 218.3 (11.0%) 350 (50%) 0.6
-2K 879.7 (44.2%) 690 (99%) 1.3
-4K 1167.6 (58.7%) 690 (99%) 1.7
-8K 1084.7 (54.5%) 690 (99%) 1.6
 For a GEMM between matrices of size $(M\times K)\times(K\times N)$, the total number of floating-point operations (FLOPs) performed is $2MKN$.
@@
 Moreover, the lower power draw of the Gaudi 2 relative to its TDP suggests that naïve TDP comparisons can be misleading, emphasizing the need for empirical evaluation.
+Table 1: Throughput and power measurements for square FP8 GEMMs with row-wise scaling.
+The ratio of measured TFLOPS and power draw relative to their peak values are included in parentheses.
+We include the TFLOPS/Watt ratio on the right.
+H100 has 1989.9 peak FP8 GEMM TFLOPS and 700W TDP.
+Gaudi 2 has 865 peak FP8 GEMM TFLOPS and 600W TDP.
+Device (M,K,N) TFLOPS Power (W) TF/W
+Gaudi 2 1K 367.9 (42.5%) 375 (63%) 1.0
+2K 586.2 (67.8%) 460 (77%) 1.3
+4K 817.1 (94.5%) 460 (77%) 1.8
+8K 741.8 (85.8%) 490 (82%) 1.5
+H100 1K 218.3 (11.0%) 350 (50%) 0.6
+2K 879.7 (44.2%) 690 (99%) 1.3
+4K 1167.6 (58.7%) 690 (99%) 1.7
+8K 1084.7 (54.5%) 690 (99%) 1.6
 4 Comparing end-to-end results
-We conduct evaluations on different FP8 features and come to the following conclusions.
-First, E4M3 consistently outperforms E5M2 in quantization accuracy across nearly all tested configurations.
-Second, stochastic rounding during quantization has minimal impact on model accuracy and, in some cases, may even be detrimental.
-Third, dynamic scaling achieves accuracy comparable to BF16 models while eliminating the need for a calibration set.
-For evaluation, we use instruction-tuned versions of Llama v3.2 1B, v3.2 3B, v3.1 8B, and v3.3 70B.
-Unless stated otherwise, models employ dynamic row-wise scaling for all linear layers, while LM head remains in BF16.
 Table 2: The “Cited” columns include results from Kim (2025) using static “FP8 (S)” scaling from the Intel Neural Compressor (INC) library using the UltraChat 200K (Ding et al., 2023) dataset for calibration.
@@
 TruthfulQA mc2 54.0% 52.2% 55.1% 54.3%
+We conduct evaluations on different FP8 features and come to the following conclusions.
+First, E4M3 consistently outperforms E5M2 in quantization accuracy across nearly all tested configurations.
+Second, stochastic rounding during quantization has minimal impact on model accuracy and, in some cases, may even be detrimental.
+Third, dynamic scaling achieves accuracy comparable to BF16 models while eliminating the need for a calibration set.
+For evaluation, we use instruction-tuned versions of Llama v3.2 1B, v3.2 3B, v3.1 8B, and v3.3 70B.
+Unless stated otherwise, models employ dynamic row-wise scaling for all linear layers, while LM head remains in BF16.
 4.1 Dynamic vs static scaling
@@
 Therefore, dynamic quantization is the preferred approach, provided that the associated throughput reduction remains within acceptable limits.
-4.2 E4M3 vs. E5M2
 Table 3: Comparison between different FP8 data types and rounding modes for MMLU CoT 5-shot performance on instruction-tuned Llama models.
@@
 E5M2 RTN 82.2%
+4.2 E4M3 vs. E5M2
 Following on the findings of Shen et al.
@@
 Furthermore, as shown in Table 8, the GEMM throughputs using E4M3 and E5M2 are comparable, making E4M3 the preferred choice for inference even considering the smaller representational capacity of E4M3 on Gaudi 2.
-Figure 2: Process and utilization characterization of the two phases of generative LLM inference.
 4.3 Stochastic rounding
@@
 These findings suggest that while stochastic rounding during quantization theoretically retains more information, its practical benefits for FP8 inference in LLMs remains inconclusive.
+Figure 2: Process and utilization characterization of the two phases of generative LLM inference.
 5 LLM inference comparison
 5.1 LLM inference phases
-Generative LLM inference consists of a compute-bound prefill phase and a memory-bound decode phase (Patel et al., 2024; Zhong et al., 2024).
-We include a diagram highlighting the differences between the two phases in Figure 2.
-In the prefill phase, the LLM generates the first token based on the provided input prompt.
-The key-value pairs are stored in the KV cache for reuse during the generation of subsequent tokens.
+Generative LLM inference comprises two distinct phases: a compute-bound prefill phase and a memory-bound decode phase (Patel et al., 2024; Zhong et al., 2024).
+Figure 2 illustrates the key differences between these phases.
+During the prefill phase, the LLM generates the first token based on the provided input prompt.
+Throughout this process, key-value (KV) pairs are stored in the KV cache for reuse in subsequent token generations.
 The prefill phase processes the entire input in parallel, utilizing matrix-matrix operations that are typically compute-bound.
-This parallelism allows for effective hardware utilization, enabling efficient input processing.
-After the first token is generated, the process transitions to the decode phase, where output tokens are generated sequentially in an autoregressive manner based on the previous tokens.
-The decode phase involves GEMV operations, inherently limiting hardware utilization compared to the compute-bound GEMMs in the prefill phase.
+This high degree of parallelism allows for effective hardware utilization, enabling efficient input processing.
+Once the first token is generated, the process transitions to the decode phase, where output tokens are generated sequentially in an autoregressive manner based on the previous tokens.
+Unlike the prefill phase, the decode phase is constrained by memory bandwidth, as it involves GEMV operations, which inherently limit hardware utilization compared to the more parallelizable GEMMs in the prefill phase.
 Modern inference frameworks such as vLLM (Kwon et al., 2023) and TensorRT-LLM (NVIDIA, 2023) partially remedy this issue by batching multiple decoding requests, improving the computational intensity of the linear layers.
@@
 In contrast, the gain for the Gaudi 2 approaches 50%.
-Figure 5: Thin GEMM MFU comparison between the Gaudi 2 and H100 for both BF16 and FP8.
-The Gaudi 2 maintains a similar MFU for BF16 and FP8 of similar shapes while there is a noticeable drop for the H100.
-The MFU differences at the same shapes are enough to provide superior TFLOPS for the Gaudi 2 over the H100 as shown in Table 4.
 To investigate the lack of performance gains in the H100, we analyze its throughput on thin matrices, identifying it as the primary bottleneck.
@@
 (64, 4096, 4096) 144.5 253.4 133.3 133.9
+Figure 5: Thin GEMM MFU comparison between the Gaudi 2 and H100 for both BF16 and FP8.
+The Gaudi 2 maintains a similar MFU for BF16 and FP8 of similar shapes while there is a noticeable drop for the H100.
+The MFU differences at the same shapes are enough to provide superior TFLOPS for the Gaudi 2 over the H100 as shown in Table 4.
 6 Conclusion

### v3

@@
-An Investigation of FP8 Across Accelerators for LLM Inference
-Abstract
-The introduction of 8-bit floating-point (FP8) computation units in modern AI accelerators has generated significant interest in FP8-based large language model (LLM) inference.
-Unlike 16-bit floating-point formats, FP8 in deep learning requires a shared scaling factor.
-Additionally, while E4M3 and E5M2 are well-defined at the individual value level, their scaling and accumulation methods remain unspecified and vary across hardware and software implementations.
-As a result, FP8 behaves more like a quantization format than a standard numeric representation.
-In this work, we provide the first comprehensive analysis of FP8 computation and acceleration on two AI accelerators: the NVIDIA H100 and Intel Gaudi 2.
-Our findings highlight that the Gaudi 2, by leveraging FP8, achieves higher throughput-to-power efficiency during LLM inference, offering valuable insights into the practical implications of FP8 adoption for datacenter-scale LLM serving.
-FP8, LLM, Quantization, GEMM, GEMV
-1 Introduction
-The FP8 format has emerged as a promising datatype due to its lower computational overhead, reduced power consumption, and improved memory efficiency compared to commonly used formats such as FP32 and BF16 (Kalamkar et al., 2019), particularly at large batch sizes.
-With the introduction of the Intel Gaudi HPU and NVIDIA Hopper GPUs, FP8 is gaining traction as a key datatype for next-generation AI workloads (DeepSeek-AI, 2024).
-While extreme low-bit quantization has been widely studied, its throughput remains inferior to conventional 16-bit floating-point formats on commercial hardware (Lin et al., 2024b).
-In contrast, FP8 has rapidly become the preferred format due to its compatibility with existing accelerators while preserving model accuracy.
-This makes it particularly well-suited for datacenter-scale LLM inference, where both high throughput and minimal accuracy degradation are critical.
-In some accelerators, FP8 can be twice as fast as BF16.
-Given the immense computational costs of training and deploying large language models (LLMs), FP8 has garnered significant attention for its potential efficiency gains.
-Despite its advantages, FP8 utilization is constrained by the widening gap between computing speed and memory bandwidth.
-A key challenge is that, while LLM sizes continue to scale, their hidden dimensions grow at a much slower rate (e.g., Llama 8B has a hidden size of 4096, but Llama v3.3 70B only increases hidden size to 8192).
-Tensor parallelism (TP) further partitions matrices into even smaller fragments, making small matrix efficiency a critical performance factor.
-Especially during the decode phase, Generalized Matrix-Vector (GEMV) and thin General Matrix-Matrix Multiply (GEMM) operations dominate, reducing Model FLOP Utilization (MFU) (Chowdhery et al., 2023; Narayanan et al., 2021).
-With the growing popularity of “reasoning” models (DeepSeek-AI, 2025), the proportion of decoding in LLM inference workloads can be expected to continue growing.
-To comprehensively address the challenges of LLM inference, we aim to quantify the various constraints that impact performance, including those introduced by FP8 quantization.
-To the best of our knowledge, we are the first to provide quantitative measurements associated with various FP8 quantization schemes.
-There is a notable lack of research providing comprehensive FP8 experimental results, particularly in studies analyzing hardware characteristics in LLM inference configurations.
-To address this gap, our work offers a detailed evaluation of FP8 performance across different hardware platforms, providing valuable insights for model developers on the trade-offs between model architecture and hardware design.
+An Inquiry into Datacenter TCO for LLM Inference with FP8
+Abstract.
+As large language models (LLMs) continue to scale, their inference demands present significant challenges, particularly due to the high power consumption of AI accelerators in datacenters.
+These facilities require specialized cooling and power management systems, substantially increasing the total cost of ownership (TCO) for cloud service providers (CSPs).
+In this work, we analyze the computational characteristics and constraints of LLM inference from a TCO perspective, focusing on two representative accelerators: the Gaudi 2 and NVIDIA H100.
+We present a generalizable framework that enables CSPs to compare and select AI accelerators according to diverse operational requirements.
+Using this model, we analyze the impact of FP8 precision and LLM inference workload characteristics as key factors influencing TCO.
+We investigate FP8 quantization, which is gaining adoption in LLM training, as a technique to improve inference throughput while maintaining cost efficiency.
+Furthermore, our analysis of LLM inference workloads reveals that performance on thin GEMMs, which dominate the decode phase, can have a greater impact than theoretical hardware peak performance.
+By studying the interaction between power consumption, quantization strategies, and hardware architecture, we offer insights that support informed deployment decisions and guide future accelerator designs to improve the TCO of LLM inference.
+LLM inference, TCO analysis, FP8 quantization
+conference: Arxiv; 2025; Arxiv
+1. Introduction
+With the rapid advance of large language models (LLMs), AI datacenters have emerged as a critical infrastructure for supporting scalable and efficient inference workloads.
+The recent emergence of ”reasoning” AI models for both proprietary models such as OpenAI’s o1 (OpenAI, 2024) and Claude 3.7 (Anthropic, 2025), as well as open-source models such as DeepSeek R1 (DeepSeek-AI, 2025), has solidified the expectation that test-time scaling of LLMs will become a central workload for cloud service providers (CSPs).
+As LLM inference requests surge in both quantity and complexity, the volume of output tokens per request is also expected to grow substantially.
+This trend places significant pressure on datacenter infrastructure, making the selection of appropriate hardware a critical factor for sustaining performance and service-level efficiency.
+In this context, competition among AI accelerator manufacturers has intensified, with each aiming to position themselves to take advantage of the market.
+Hardware selection has a direct impact on the total cost of ownership (TCO), which in turn influences the overall profitability and operational viability of AI datacenters.
+This necessitates a detailed examination of the various components that impact TCO, ultimately aiming to identify effective strategies for minimizing it.
+To provide practical guidance for CSPs, we construct an analytical model to compare hardware options, taking into account both facility costs and target performance levels.
+Although extensive research has explored the design of advanced AI accelerators (Micikevicius et al., 2022; Lee et al., 2025) and LLM quantization techniques (Lin et al., 2024; Ashkboos et al., 2024; Xiao et al., 2024), there remains a lack of rigorous investigation into how these developments impact the TCO for datacenter-scale LLM inference.
+Building upon our model, we analyze several key factors that influence TCO and discuss potential approaches for its optimization.
+First, we analyze the impact of data precision in LLMs, with a particular focus on FP8 quantization, which has seen increasing adoption in recent state-of-the-art models such as DeepSeek v3 (DeepSeek-AI, 2024) and Llama 4 (Meta, 2025).
+Quantization offers the additional benefit that modern AI accelerators often include dedicated hardware units designed to accelerate low-precision computation, thereby improving throughput.
+In particular, FP8 has emerged as a promising datatype due to its lower computational overhead, reduced power consumption, and improved memory efficiency compared to commonly used formats such as FP32 and BrainFloat16 (BF16) (Kalamkar et al., 2019), especially for large batch sizes.
+With the adoption of FP8 by next-generation accelerators such as Intel’s Gaudi HPU and NVIDIA’s Hopper GPU, this format is gaining traction in cutting-edge LLMs.
+In this context, we explore how different FP8 quantization strategies impact both accuracy and throughput, examining their effects at two levels: matrix multiplication operations and end-to-end LLM inference.
+Second, we investigate the behavior of LLM inference workloads when executed with FP8 quantization, aiming to understand its impact on throughput and resource utilization.
+To leverage quantization in a cost-effective manner, it is essential to understand the computational characteristics of LLM inference, which differ significantly from those of training.
+Even within the inference process, throughput varies across two distinct phases: the prefill phase and the decode phase.
+In particular, the decoding phase is dominated by Generalized Matrix-Vector (GEMV) operations and thin General Matrix-Matrix Multiplication (GEMM), both of which exhibit low computational intensity (CI) and lead to reduced Model FLOP Utilization (MFU) (Chowdhery et al., 2023; Narayanan et al., 2021).
+Furthermore, performance varies significantly across hardware platforms, depending on factors such as memory interface bandwidth and matrix computation unit throughput.
+Therefore, we conduct an in-depth analysis of the characteristics of the H100 and Gaudi2 in the context of LLM inference workloads.
+We further identify the performance limitations of each accelerator by examining their inherent architectural bottlenecks.
 To summarize our contributions:
-Figure 1: TFLOPS (top row) and TFLOPS/Watt (bottom row) comparison between the Gaudi 2 and H100 using FP8 in the decode phase with batch size set to 64.
-We measure inference throughput using Optimum Habana v1.15.0 for the Gaudi 2 and TensorRT-LLM v0.16.0 for the H100.
-For short sequences, the Gaudi 2 has an advantage both in absolute throughput and in TFLOPS/Watt.
-At longer sequences, the H100 achieves higher TFLOPS due to its superior memory bandwidth, but is less energy efficient due to higher power consumption.
-  1. 1.
-    We analyze scaling factor configurations for FP8 formats and quantify key LLM inference constraints in datacenter environments, providing a unified framework to compare throughputs across different inference stages, model sizes, and sequence lengths.
-  2. 2.
-    We analyze the computational characteristics of LLM inference on the Gaudi 2 and H100 across both the prefill and decode phases.
-In the prefill phase, where compute performance is the key factor, H100’s throughput is 50% higher than Gaudi 2’s.
-However, since the decode phase dominates overall inference time, accelerator utilization becomes more critical, and Gaudi 2 achieves 50% higher throughput in this phase.
-This highlights the differing performance trade-offs between the two accelerators across LLM inference workloads.
-2 Background
-With the explosive growth of LLM sizes in recent years, numerous techniques specific to generative LLMs have emerged to reduce memory requirements and improve output speeds.
-In this work, we focus on the large-batch serving environment often found in data centers, where multiple requests for LLM inference run simultaneously under power constraints at both the rack and data center levels.
-Various methods have been proposed to improve LLM serving efficiency, with quantization being one of the most widely studied.
-However, a key challenge in LLM quantization is the presence of activation outliers.
-Transformer models often contain large outlier values in their activations, which, if not properly addressed during quantization, can lead to substantial quality degradation (Sun et al., 2024a).
-To mitigate this issue, several quantization techniques have been developed, aiming to improve robustness and accuracy while maintaining computational efficiency.
-2.1 Weight-only quantization
-Due to the quantization difficulties posed by activation outliers, many studies have proposed several weight-only quantization (WoQ) methods(Lin et al., 2024a; Frantar et al., 2023; Lee et al., 2023), while retaining activations in BF16 or FP16 format.
-However, these approaches often suffer from low throughput at larger batch sizes.
-To enable computations, quantized weights must be converted into the same data type as activations (e.g.
-BF16), requiring online dequantization.
-This is handled by CUDA cores on NVIDIA GPUs and TPC cores on Gaudi HPUs, which operate an order of magnitude slower than their respective matrix multiplication units.
-Therefore, dequantization frequently becomes a bottleneck even at moderate batch sizes.
-Also, because GEMM computations are conducted at higher precisions, WoQ methods are fundamentally limited to having the same throughput as unquantized GEMMs at best.
-Even when high throughput is achieved, kernel implementations often need to be re-implemented for different devices, including those from the same vendor.
-For instance, the Marlin kernel (Frantar et al., 2024), optimized for int4-weight and float16-activations on the A100, is suboptimal for the H100.
-As a result, the Machete kernel (Wilkinson, 2024) was developed specifically to leverage the new features of the Hopper architecture, employing different design trade-offs.
-These challenges arise due to variations in cache sizes, hardware features, and compute-to-memory bandwidth ratios, all of which necessitate different optimizations to achieve high performance.
-Consequently, adapting WoQ methods across different accelerators imposes significant implementation overhead and increases development costs.
-2.2 INT8 quantization
-The INT8 quantization format described in Xiao et al.
-(2024) is closely related to our work, requiring quantization of both weights and activations to leverage INT8 tensor cores to accelerate computations.
-While INT8 quantization has been widely adopted in fields such as computer vision, its limited dynamic range presents challenges in achieving full accuracy for LLM inference.
-FP8 thus has an advantage over INT8 as integer representations are inherently less robust to outliers compared to floating-point numbers (Shen et al., 2024; Kurtic et al., 2024).
-2.3 FP8 quantization
-The FP8 format can be categorized into two variants based on the allocation of bits to the exponent and mantissa: the E4M3 format and the E5M2 format.
-E4M3 offers higher precision compared to the E5M2 format, which instead has a higher dynamic range.
-To address the smaller dynamic range of FP8 compared to BF16, scaling factors are employed to adjust the inputs before casting to FP8.
-This ensures that the values fit within the FP8 range, maximizing the utilization of its limited dynamic range.
-Per-tensor scaling, which computes a single scaling factor for each matrix.
-For more granular control, per-token scaling can be applied to activations, while per-channel scaling is used for weights, enabling finer adjustments.
-Group-wise scaling factors (DeepSeek-AI, 2024) have also been proposed.
-2.4 Scaling factor computation
-A critical consideration for activation quantization is the overhead of converting activations to lower precision.
-While non-linear operations require 16-bit or even 32-bit precision, activations must be quantized online immediately before GEMM computations utilize low-precision compute units.
-This introduces non-trivial overhead, limiting the throughput gains from reduced precision.
-Activation quantization also requires estimating input statistics to determine appropriate scaling factors.
-A direct approach, measuring statistics for each activation before quantization, would necessitate two memory accesses: one for collecting statistics and another for quantization, incurring high latency.
-To mitigate this, many libraries use a calibration set to precompute scaling factors based on a representative sample of activations.
-A key limitation of static scaling is that it typically enforces per-tensor scaling of activations, as token-wise variations are unknown in advance.
-For instance, altering the order of sequences during batched decoding would shift token positions.
-Moreover, using position-dependent scaling factors would require fixed input sizes.
-While per-channel scaling (Xiao et al., 2024) is possible, the same scaling factors are still applied to all tokens, potentially reducing accuracy.
-In contrast, dynamic scaling is more efficient when applied separately to each token as measurement and quantization can be combined in a single memory access.
-It also enables group-wise quantization, as implemented in DeepSeek-AI (2024) with a group size of 128, improving output quality through finer-grained scaling factors.
-A potential optimization for both scaling methods is to merge quantization with other memory-bound operations.
-In Pre-LN Transformers (Xiong et al., 2020), activation quantization can be fused with RMSNorm.
-Similarly, element-wise multiplication in SwiGLU (Shazeer, 2020) can also be fused.
-Frameworks such as TensorRT-LLM (NVIDIA, 2023) v0.16.0 have already implemented these optimizations for FP8 activation quantization.
-They have yet to be implemented for Gaudi HPUs as of the time of writing.
-2.5 LLM evaluation
-LLM evaluation has grown in importance as new LLMs compete for prominence.
-Traditionally, WikiText-2 perplexity (Merity et al., 2017) has been widely used due to its ease of measurement and relationship with the training loss.
-More recently, Massive Multitask Language Understanding (MMLU) (Hendrycks et al., 2021) has emerged as the de facto standard to assess LLM quality, evaluating knowledge in a wide range of domains.
-However, multiple-choice benchmarks may fail to capture the full impact of quantization.
-For instance, the Llama v3 405B model (Llama Team, 2024) applies max-absolute (maxabs) clipping when calculating scaling factors despite there being no effect in standard evaluation metrics because the model may generate corrupted outputs for high-perplexity tokens, such as dates, without maxabs clipping.
-Moreover, evaluation should extend beyond pre-trained base models to instruction-tuned models that have undergone alignment.
-If quantization is applied after pre-training but before alignment, the alignment process must be quantization-aware, requiring each quantization algorithm to be integrated with each alignment method while maintaining alignment quality, significantly increasing complexity.
-For this reason, we conduct all evaluations on instruction-tuned models, where quantization effects are more representative of real-world deployment scenarios.
-3 Matrix multiplication comparison
-3.1 Differences in hardware capabilities
+  1. (1)
+    We present a scalable TCO estimation model that serves as a generalizable framework for selecting optimal hardware configurations tailored to the needs of CSPs.
+While previous studies (Patel et al., 2024; Griggs et al., 2024) have analyzed throughput-cost trade-offs, they are often limited to specific deployment scenarios.
+In contrast, our model accommodates a wide range of cost structures and performance requirements, enabling flexible evaluation across different CSP contexts.
+Moreover, the model offers intuitive guidance for TCO optimization from a throughput-centric perspective.
+  2. (2)
+    Using this model, we investigate two key factors influencing TCO: FP8 numerical precision and LLM inference workload characteristics.
+Despite FP8’s growing popularity as a low-precision format that offers favorable trade-offs between throughput and memory efficiency, prior work has largely overlooked its empirical evaluation, particularly in the context of hardware-specific behavior during LLM inference.
+To address this gap, we conduct a comprehensive analysis of FP8 performance across different accelerator platforms.
+In addition, we provide a focused analysis of LLM decoding throughput that goes beyond a simple inspection of hardware specifications.
+Under FP8-quantized conditions, we further explore the performance characteristics and bottlenecks of each device within LLM workloads, offering insights that inform both hardware design and deployment strategies.
+2. Exploring the TCO-throughput trade-off
+People who buy and sell chips think about the cost of chips.
+People who operate data centers think about the cost of operations.
+… Our TCO is so good that even when the competitor’s chips are free, it’s not cheap enough.
+Jensen Huang, Founder & CEO of NVIDIA
+2.1. Understanding the TCO model
+Figure 1.
+TCO comparison table for fixed traffic.
+For any desired task, given the throughput ratio between accelerators A and B $R_{Th}$ and server price ratio $R_{SC}$, the table shows the TCO of accelerator A as a proportion of the TCO of accelerator B, with $C_{S}=C_{I}$ assumed for clarity.
+A is therefore more cost-efficient than B when the obtained value is less than 1.
+The lower the value, the greater the cost-effectiveness of A.
+When calculating the costs involved in running AI accelerators for datacenter operators, one must note that the purchase price of the chips themselves, while significant, constitute only part of the Total Cost of Ownership (TCO).
+For example, the purchase cost of the networking, cooling, and power infrastructure necessary to maintain the processors is frequently as large as the purchase cost of the chips and may even exceed it (Cui et al., 2017).
+As AI inference workloads constitute an ever greater portion of datacenter operations, achieving low cost and high efficiency requires careful planning.
+The choice of computing systems impacts both the throughput and the power consumption, both of which have a substantial influence on TCO.
+Consider a scenario in which a CSP must choose between accelerators for a datacenter.
+In this scenario, two alternatives exist: system A, which is lighter-weight, delivers lower peak specifications, and has a lower purchase price, and system B, which provides substantially higher peak specifications but comes with a higher per accelerator purchase price.
+To identify which hardware option yields higher profitability, we must construct a model to compare TCO between the two hardware options.
+When purchasing a server, it contains not only the chip but also networking, storage, and other associated components required for the server to function.
+Also, mulltiple chips may be located in a single server, whose costs must also be accounted for.
+In the following discussion, we denote $ServerCost$ as the term that combines the purchase costs of the multiple accelerator chips, CPUs, storage, and networking devices required to constitute a operational server.
+$InfraCost$ indicates the infrastructure costs associated with operating the datacenter, including cooling and power provision, divided by the number of accelerators.
+We simplify the analysis by including operating expenses (OpEx), such as electricity consumption, in $InfraCost$.
+To understand the infrastructure cost, note that while the cost of electricity increases almost linearly with regard to the number of processors, the cost of the associated infrastructure, such as power racks and cooling equipment, is nearly constant.
+Because datacenter racks can only provide a limited amount of power, the per-chip cost of infrastructure is inversely proportional to the number of servers that can fit in a single rack.
+Therefore, the benefits of lower power consumption are twofold: the electricity consumption goes down and the number of processors per rack increases, spreading fixed infrastructure costs among a greater number of servers.
+For most datacenters, the cost of electricity per se is outweighed the cost of the rack and other equipment, even when amortized over several years, leading $InfraCost$ to be most heavily affected by the number of chips per rack.
+We denote $R_{SC}$, $R_{IC}$, and $R_{Th}$ as the relative server costs, per server infrastructure costs, and throughputs between a single server of computing systems A and B, respectively, and define them as follows:
+$R_{SC}=\frac{ServerCost_{A}}{ServerCost_{B}},\quad R_{IC}=\frac{InfraCost_{A}}{InfraCost_{B}},\quad R_{Th}=\frac{Throughput_{A}}{Throughput_{B}}$
+Regardless of the devices in use, the amount of traffic remains constant.
+We therefore make an iso-traffic assumption and use $N$ to represent the number of B servers required to process the expected traffic.
+We use $C_{S}=ServerCost_{B}$ and $C_{I}=InfraCost_{B}$ to represent the cost components attributed to the server and the infrastructure cost per server when using system B.
+Making the assumption that the cost of infrastructure is the same for systems A and B, i.e., $R_{IC}=1$, derive the following equation:
+(1) $\frac{TCO_{A}}{TCO_{B}}=\frac{C_{S}\times R_{SC}\times\frac{N}{R_{Th}}+C_{I}\times\frac{N}{R_{Th}}}{C_{S}\times N+C_{I}\times N}$
+The numerator captures the TCO incurred when using the lower-performance system A to meet the same throughput target as system B, while the denominator reflects the baseline TCO of using system B for a given amount of traffic.
+As total traffic remains constant regardless of the underlying device, purchasing system A will require $R_{Th}/N$ nodes, leading to possibly increased infrastructure costs, depending on the throughput ratio $R_{Th}$.
+At the same time, its lower per-chip cost, reflected in the server price ratio $R_{SC}$, could result in a lower TCO.
+Based on this model, Figure 1 illustrates the TCO comparison between two computing systems, expressed as an $A/B$ ratio.
+The x-axis represents the server price ratio, and the y-axis represents the per server throughput ratio of the two systems for the desired task.
+Since exact costs are often confidential in real-world data centers, expressing TCO in relative ratios allows for flexible adaptation to various operational contexts.
+In deriving the values in Figure 1, we assumed $C_{S}=C_{I}$, assuming that server and infrastructure costs contribute equally (i.e., 50% each) to the overall TCO, thereby allowing these terms to cancel out across the numerator and denominator.
+We also assume that both systems consume the same power per server and that the same number of servers fit in a single rack, i.e., $R_{IC}=1$.
+In practice, however, it may be appropriate to substitute variables to reflect each specific deployment scenario.
+The values in Figure 1 are an example of the TCO ratio between two hardware options.
+The lower the value, the less expensive, and hence more cost-effective, system A becomes and the more appealing it is to adopt.
+Conversely, a ratio above 1.0 suggests that system B is preferable from a TCO perspective.
+2.2. TCO factor analysis
+The TCO analysis presented so far offers a flexible and adaptable methodology for evaluating system options across diverse data center environments.
+In the implemented TCO model, both the server price ratio and throughput ratio can vary significantly depending on factors such as server acquisition costs, accuracy constraints, and the specific tasks being executed.
+The cost of processors may change depending on negotiations with manufacturers and are generally confidential.
+A key contribution of the analysis is that it can incorporate throughput ratios of any desired task, e.g., LLM decoding as opposed to LLM training or prefill, allowing users to take the intricacies of hardware-software co-optimization into account.
+For instance, the service may require that only models that pass a quality assurance (QA) review are allowed, and that the model meets specific latency requirements when deployed.
+All of these factors can be incorporated into the TCO model to provide insights.
+For example, the prefill and decode phases of LLM inference exhibit markedly different throughput profiles.
+As shown in Patel et al.
+(2024), it is possible to use different chips and configurations for the prefill and decode to optimize TCO as their relative throughputs for the prefill and decode tasks are different.
+When these different throughputs are compared using our TCO model, it becomes easier to grasp why it is profitable to utilize different hardware and configurations for different inference phases.
+To build an efficient datacenter environment, CSPs must consider complex trade-offs between functionality, performance, and cost.
+Quantization has emerged as a viable solution to overcome some of these limitations by reducing memory usage.
+Leveraging the inherent error resilience of LLMs enables reduced bit-width representations of model parameters while maintaining comparable output quality.
+Consequently, incorporating these lower-precision modes into TCO analysis becomes essential.
+In the following section, we explore FP8 quantization as a particularly important technique.
+3. FP8 for high-throughput LLM inference
+In this section, we provide an overview of FP8 as a numerical format, analyze how it is implemented across hardware backends, and evaluate its performance in GEMM operations, one of the core building blocks of LLM inference.
+3.1. Motivation for FP8 precision
+While BF16 precision is commonly used in recent LLMs, there is substantial opportunity to boost throughput and to minimize TCO by operating at a lower precision level, FP8, within acceptable bounds of accuracy degradation.
+Especially with the emergence of next-generation foundation models such as DeepSeek v3 and LLaMA 4 whose parameters were trained in FP8 precision, the FP8 GEMM has become increasingly central to achieving high throughput at scale.
+Moreover, even for models originally trained using BF16, post-training quantization (PTQ) techniques now enable effective conversion to FP8 without significant accuracy degradation.
+However, the effectiveness of low-precision computation is highly dependent on the underlying hardware architecture.
+As precision decreases, achieving full hardware utilization becomes more challenging, directly affecting throughput.
+Therefore, hardware-specific characteristics are critical in determining actual performance at lower precision levels.
+In this section, we explore the characteristics and implications of Llama 4 style FP8 precision, analyzing its computational behavior and potential for performance optimization.
+3.2. Differences in hardware capabilities
 We compare implementations of FP8 GEMM between NVIDIA GPUs and Intel Gaudi HPUs, identifying key differences despite both adhering to the FP8 specification (Micikevicius et al., 2022).
@@
 Software optimizations, such as applying FP32 accumulation to only one in four WGMMA instructions, reduce error but increase kernel complexity and remain less precise than full FP32 accumulation.
-In contrast, Gaudi HPUs always accumulate in FP32, ensuring higher numerical precision.
+In contrast, Gaudi HPUs always accumulate in FP32 (Lee et al., 2025), ensuring higher numerical precision.
 (E4M3 range) The Gaudi 2 follows the IEEE specification for special values, unlike NVIDIA GPUs, which use a single special value representation (Noune et al., 2022).
@@
 This has been addressed in the Gaudi 3, but we were unable to test these in our experiments.
-(Power-of-2 scaling) Gaudi HPUs allow modification of floating-point exponent biases to accelerate scaling factor application.
-The Gaudi 2 supports fixed hardware-accelerated scaling factors of $2^{-8},2^{-4},2^{0},2^{4}$ for E4M3 while the Gaudi 3 extends this to arbitrary powers of 2 between $2^{-32}$ and $2^{31}$.
+(Power-of-2 scaling) Gaudi HPUs allow the modification of floating-point exponent biases to accelerate scaling factor application.
+The Gaudi 2 supports fixed hardware-accelerated scaling factors of $2^{-8},2^{-4},2^{0},2^{4}$ for E4M3, while the Gaudi 3 extends this to arbitrary powers of 2 between $2^{-32}$ and $2^{31}$.
 However, this feature is limited to GEMMs with per-tensor scaling factors.
@@
 (Sparsity) NVIDIA GPUs support semi-structured sparsity acceleration, potentially doubling tensor core peak throughput.
-However, despite attempts to leverage sparsity in LLMs (Frantar & Alistarh, 2023; Sun et al., 2024b), dense GEMMs remain dominant due to accuracy loss and limited speedups.
+However, despite attempts to leverage sparsity in LLMs (Frantar and Alistarh, 2023; Sun et al., 2024), dense GEMMs remain dominant due to accuracy loss and limited speedups.
 Gaudi HPUs do not support sparsity acceleration.
-3.2 GEMM throughput measurements
-For a GEMM between matrices of size $(M\times K)\times(K\times N)$, the total number of floating-point operations (FLOPs) performed is $2MKN$.
-This is derived from the $M\times N$ dot products of length $K$, where each element undergoes one multiplication and one addition.
-Following the convention where FLOPS denotes FLOPs per second, we compute throughput in FLOPS using the theoretical FLOPs and the measured latency.
-We can thus calculate the MFU by dividing observed throughput by peak throughput.
-As LLM computations are dominated by matrix multiplications, GEMM throughput serves as an upper bound for model end-to-end performance.
-While accelerator specifications list peak throughput values, actual throughput rarely reaches this limit, particularly for small matrices.
-Table 1 presents GEMM throughput and power consumption measurements for row-wise scaled FP8 GEMM.
-We conducted measurements on NVIDIA H100 GPUs using the NGC PyTorch 24.12 image and on Intel Gaudi 2 HPUs using the Synapse AI v1.19.0 image.
-Power consumption was recorded via the ”nvidia-smi” and ”hl-smi” utilities, respectively.
-Input casting overheads were excluded from the measurements.
-Additional results for GEMM throughputs under different conditions are provided in Tables 8 and 9.
-Our results show that the Gaudi 2 achieves higher utilization than the H100, particularly for small matrices.
-The performance drop for smaller matrices is steeper on the H100, with the Gaudi 2 providing a higher TFLOPS at matrix sizes of 1K.
-Also, the Gaudi 2 exhibits lower power consumption than its stated 600W TDP, whereas the H100 approaches its peak 700W TDP even at moderate utilizations.
-Unlike NVIDIA GPUs, which incorporate multiple small matrix multiplication accelerators, Gaudi HPUs feature a few large matrix accelerators.
-As a result, Gaudi HPUs requires fewer input elements per cycle to fully utilize compute resources, thereby reducing first-level memory bandwidth requirements and improving efficiency (Intel, 2022, 2024).
-Additionally, Gaudi HPUs leverage a graph compiler to dynamically reconfigure the MME size based on the target GEMM dimensions, optimizing resource utilization.
-This superior efficiency of the Gaudi 2 for small matrices is particularly relevant given that LLM hidden dimensions typically range between 1K (e.g., Llama 1B) and 8K (e.g., Llama 70B), with tensor parallelism further reducing matrix sizes.
-Moreover, the lower power draw of the Gaudi 2 relative to its TDP suggests that naïve TDP comparisons can be misleading, emphasizing the need for empirical evaluation.
-Table 1: Throughput and power measurements for square FP8 GEMMs with row-wise scaling.
-The ratio of measured TFLOPS and power draw relative to their peak values are included in parentheses.
-We include the TFLOPS/Watt ratio on the right.
-H100 has 1989.9 peak FP8 GEMM TFLOPS and 700W TDP.
-Gaudi 2 has 865 peak FP8 GEMM TFLOPS and 600W TDP.
-Device (M,K,N) TFLOPS Power (W) TF/W
+Table 1.
+Throughput and power measurements for square FP8 GEMMs with row-wise scaling.
+Ratios of measured TFLOPS and power to their peak values are shown in parentheses.
+H100 has peak FP8 throughput of 1989.9 TFLOPS with a TDP of 700W, and Gaudi 2 has 865 TFLOPS with a TDP of 600W.
+Device (M,K,N) TFLOPS Power (W) TFLOPS/W
 Gaudi 2 1K 367.9 (42.5%) 375 (63%) 1.0
@@
 8K 1084.7 (54.5%) 690 (99%) 1.6
-4 Comparing end-to-end results
-Table 2: The “Cited” columns include results from Kim (2025) using static “FP8 (S)” scaling from the Intel Neural Compressor (INC) library using the UltraChat 200K (Ding et al., 2023) dataset for calibration.
+3.3. GEMM throughput measurements
+For a GEMM between matrices of size $(M\times K)\times(K\times N)$, the total number of floating-point operations (FLOPs) performed is $2MKN$.
+This is derived from the $M\times N$ dot products of length $K$, where each element undergoes one multiplication and one addition.
+Following the convention where FLOPS denotes FLOPs per second, we compute throughput in FLOPS using the theoretical FLOPs and the measured latency.
+We can thus calculate the MFU by dividing observed throughput by peak throughput.
+As LLM computations are dominated by matrix multiplications, GEMM throughput serves as an upper bound for model end-to-end performance.
+While accelerator specifications list peak throughput values, actual throughput rarely reaches this limit, particularly for small matrices.
+Table 1 presents GEMM throughput and power consumption measurements for row-wise scaled FP8 GEMM.
+We conducted measurements on NVIDIA H100 GPUs using the NGC PyTorch 25.03 image and on Intel Gaudi 2 HPUs using the Synapse AI v1.12.0 image.
+Power consumption was recorded via the “nvidia-smi” and “hl-smi” utilities, respectively.
+Input casting overheads were excluded from the measurements.
+Additional results for GEMM throughputs under different conditions are provided in Tables 2 and 3.
+Our results show that the Gaudi 2 achieves higher utilization than the H100, particularly for small matrices.
+The performance drop for smaller matrices is steeper on the H100, with the Gaudi 2 providing a higher TFLOPS at matrix sizes of 1K.
+Also, the Gaudi 2 exhibits lower power consumption than its stated 600W TDP, whereas the H100 approaches its peak 700W TDP even at moderate utilizations.
+Unlike NVIDIA GPUs, which incorporate multiple small matrix multiplication accelerators, Gaudi HPUs feature a few large matrix accelerators.
+As a result, Gaudi HPUs requires fewer input elements per cycle to fully utilize compute resources, thereby reducing first-level memory bandwidth requirements and improving efficiency (Intel, 2022, 2024).
+Additionally, Gaudi HPUs leverage a graph compiler to dynamically reconfigure the MME size based on the target GEMM dimensions, optimizing resource utilization.
+This superior efficiency of the Gaudi 2 for small matrices is particularly relevant given that LLM hidden dimensions typically range between 1K (e.g., Llama 1B) and 8K (e.g., Llama 70B), with tensor parallelism further reducing matrix sizes.
+Moreover, the lower power draw of the Gaudi 2 relative to its TDP, as shown in Table 1, suggests that naïve TDP comparisons can be misleading, emphasizing the need for empirical evaluation.
+Table 2.
+Gaudi 2 throughput in TFLOPS for scaled FP8 GEMM for square matrices of the given size, excluding quantization overhead.
+Measured throughput relative to the peak FP8 throughput (865 TFLOPS) is included in parentheses.
+Hardware acceleration is only available for per-tensor scaling.
+Type Size Per-Row Per-Tensor HW Accel.
+E4M3 1K 494 (57.1%) 494 (57.1%) 494 (57.1%)
+2K 506 (58.5%) 641 (74.1%) 641 (74.2%)
+4K 735 (84.9%) 796 (92.1%) 801 (92.6%)
+8K 742 (85.7%) 822 (95.0%) 852 (98.4%)
+E5M2 1K 306 (35.4%) 494 (57.1%) 493 (57.0%)
+2K 506 (58.5%) 642 (74.2%) 642 (74.2%)
+4K 735 (84.9%) 802 (92.7%) 802 (92.7%)
+8K 726 (83.9%) 825 (95.4%) 825 (95.4%)
+Table 3.
+H100 throughput (TFLOPS) for scaled FP8 GEMMs on square matrices of the given size, excluding quantization overhead.
+Ratios relative to peak FP8 throughput are shown in parentheses.
+Only E4M3 results are reported, as Python functions for E5M2 GEMM are unavailable for GPUs.
+Accum. Size Per-Row Per-Tensor
+FP32 1K 217 (11.0%) 186   (9.4%)
+2K 299 (15.1%) 840 (42.4%)
+4K 362 (18.3%) 1099 (55.5%)
+8K 396 (20.0%) 1300 (65.7%)
+Fast 1K 237 (12.0%) 147   (7.4%)
+2K 810 (40.9%) 896 (45.3%)
+4K 1136 (57.4%) 1205 (60.9%)
+8K 1123 (56.8%) 1388 (70.1%)
+4. FP8 configuration impact on accuracy
+Table 4.
+The “Cited” columns include results from Kim (2025) using static “FP8 (S)” scaling from the Intel Neural Compressor (INC) library using the UltraChat 200K (Ding et al., 2023) dataset for calibration.
 The “Measured” column includes results for dynamic “FP8 (D)” scaling.
-Reference BF16 accuracies are included for both columns to control for evaluation condition differences.
+BF16 accuracies are included for both columns to control for evaluation condition differences.
 Llama v3.1 8B Inst. Cited Measured
@@
 TruthfulQA mc2 54.0% 52.2% 55.1% 54.3%
-We conduct evaluations on different FP8 features and come to the following conclusions.
-First, E4M3 consistently outperforms E5M2 in quantization accuracy across nearly all tested configurations.
-Second, stochastic rounding during quantization has minimal impact on model accuracy and, in some cases, may even be detrimental.
-Third, dynamic scaling achieves accuracy comparable to BF16 models while eliminating the need for a calibration set.
-For evaluation, we use instruction-tuned versions of Llama v3.2 1B, v3.2 3B, v3.1 8B, and v3.3 70B.
-Unless stated otherwise, models employ dynamic row-wise scaling for all linear layers, while LM head remains in BF16.
-4.1 Dynamic vs static scaling
-Section 2 describes the trade-offs between static and dynamic scaling for activation quantization.
+In this section, we investigate how different FP8 quantization configurations affect the accuracy of large language models.
+Specifically, we evaluate the impact of scaling strategies (static vs.
+dynamic), format types (E4M3 vs.
+E5M2), and rounding methods (stochastic vs.
+deterministic).
+Our experiments are conducted on instruction-tuned variants of Llama v3.2 1B, v3.2 3B, v3.1 8B, and v3.3 70B.
+Unless otherwise specified, dynamic row-wise scaling is applied to all linear layers, while the LM head remains in BF16.
+4.1. Dynamic vs static scaling
+We examine the trade-offs between static and dynamic scaling strategies for activation quantization.
 Static scaling offers higher throughput for large matrices as scaling factors do not need to be computed for each input.
@@
 In contrast, dynamic scaling enhances output quality by assigning separate scaling factors to each token.
-Table 2 presents a comparative analysis of model output quality on static vs.
-dynamic FP8 scaling.
-The results indicate that dynamic quantization obtains results similar to the original BF16 models, whereas static quantization leads to noticeable accuracy degradation.
-While various techniques, such as those proposed by Xiao et al.
-(2024); Liu et al.
-(2024); Ashkboos et al.
-(2024), can potentially mitigate this degradation, they still pose a risk as the scaling factors obtained during calibration may be suboptimal for unseen inputs.
-Therefore, dynamic quantization is the preferred approach, provided that the associated throughput reduction remains within acceptable limits.
-Table 3: Comparison between different FP8 data types and rounding modes for MMLU CoT 5-shot performance on instruction-tuned Llama models.
+Native FP8 foundation models such as DeepSeek v3 and Llama 4 use dynamic scaling for both training and inference.
+Table 4 presents a comparative analysis of model output quality under static and dynamic FP8 scaling.
+The results show that dynamic quantization achieves accuracy comparable to the original BF16 models, while static quantization often results in noticeable degradation.
+Although various techniques, such as smoothing (Xiao et al., 2024) and rotation-based methods (Liu et al., 2024; Ashkboos et al., 2024), can help mitigate this degradation, they rely on calibration data and may yield suboptimal scaling factors for unseen inputs.
+As a result, dynamic quantization remains the preferred approach, assuming the associated throughput overhead is within acceptable bounds.
+Table 5.
+Comparison between different FP8 data types and rounding modes for MMLU CoT 5-shot performance on instruction-tuned Llama models.
 Stochastic rounding provides little or no benefit to accuracy while E5M2 is detrimental, especially for smaller models.
@@
 E5M2 RTN 82.2%
-4.2 E4M3 vs. E5M2
-Following on the findings of Shen et al.
-(2024), which demonstrated that the E4M3 format achieves superior accuracy on language tasks compared to E5M2, we extend this analysis to instruction-tuned models in Table 3.
-To determine the optimal format for inference, we conducted a comparative evaluation of E4M3 and E5M2, measuring MMLU accuracy using LM Evaluation Harness (v0.4.7, (Gao et al., 2024)).
-Our experimental results consistently indicate that E4M3 outperforms E5M2 across all evaluated scenarios.
-Furthermore, as shown in Table 8, the GEMM throughputs using E4M3 and E5M2 are comparable, making E4M3 the preferred choice for inference even considering the smaller representational capacity of E4M3 on Gaudi 2.
-4.3 Stochastic rounding
-Hardware-accelerated stochastic rounding during FP8 quantization is a distinctive feature of Gaudi HPUs, similar to the approach proposed by Fan et al.
-(2024), and is unavailable in NVIDIA GPUs.
+4.2. E4M3 vs. E5M2
+Prior work has shown that E4M3 yields better accuracy than E5M2 on language tasks (Shen et al., 2024).
+We extend this analysis to instruction-tuned models, as summarized in Table 5.
+To identify the optimal format for inference, we compare the two using MMLU accuracy measured with LM Evaluation Harness (v0.4.7 (Gao et al., 2024)).
+Across all evaluated models, E4M3 consistently outperforms E5M2.
+Moreover, Table 2 shows that GEMM throughputs for both formats are comparable, making E4M3 the preferred choice despite its smaller representational range on the Gaudi 2.
+Figure 2. Process and utilization characterization of the two phases of generative LLM inference.
+4.3. Stochastic rounding
+Hardware-accelerated stochastic rounding for FP8 quantization is a unique feature of Gaudi HPUs, inspired by techniques such as those proposed in (Fan et al., 2024), and is not supported on NVIDIA GPUs.
 Equation 2 formalizes the concept, where a higher-precision value $x$ is rounded up to $x_{up}$ or down to $x_{down}$ stochastically based on the distance from $x$.
-$x_{quantized}=\begin{cases}x_{up}\quad(p=\frac{x-x_{down}}{x_{up}-x_{down}})\\ x_{down}\quad(p=\frac{x_{up}-x}{x_{up}-x_{down}})\end{cases}$ (1)
+(2) $x_{quantized}=\begin{cases}x_{up}\quad(p=\frac{x-x_{down}}{x_{up}-x_{down}})\\ x_{down}\quad(p=\frac{x_{up}-x}{x_{up}-x_{down}}).\end{cases}$
 Stochastic rounding is expected to preserve more of the original numerical information post-quantization, potentially leading to improved model accuracy.
 However, empirical results in Table 5 indicate that stochastic rounding during quantization does not significantly enhance model accuracy.
-Furthermore, as shown in Table 6, it may even lead to accuracy degradation in certain cases.
-These findings suggest that while stochastic rounding during quantization theoretically retains more information, its practical benefits for FP8 inference in LLMs remains inconclusive.
-Figure 2: Process and utilization characterization of the two phases of generative LLM inference.
-5 LLM inference comparison
-5.1 LLM inference phases
+Furthermore, it may even lead to accuracy degradation in certain cases (Fan et al., 2024).
+These findings suggest that while stochastic rounding during quantization theoretically retains more information, its practical benefits for FP8 inference in LLMs remain inconclusive.
+5. Phase-aware analysis of FP8 LLM inference
+LLM inference comprises two distinct computational phases, prefill and decode, each exhibiting different performance characteristics and hardware bottlenecks.
+In this section, we analyze how FP8 quantization affects each phase, focusing on the interaction between matrix shapes, computational intensity, and hardware utilization.
+Figure 3.
+Comparison of measured throughput and the roofline throughputs obtained from hardware specifications.
+While the H100 offers significantly higher theoretical throughput, Gaudi 2 achieves superior measured performance in many decoding settings.
+Measurements were conducted using a 400W power cap per chip for fair comparison.
+5.1. LLM inference phases
 Generative LLM inference comprises two distinct phases: a compute-bound prefill phase and a memory-bound decode phase (Patel et al., 2024; Zhong et al., 2024).
@@
 Once the first token is generated, the process transitions to the decode phase, where output tokens are generated sequentially in an autoregressive manner based on the previous tokens.
-Unlike the prefill phase, the decode phase is constrained by memory bandwidth, as it involves GEMV operations, which inherently limit hardware utilization compared to the more parallelizable GEMMs in the prefill phase.
+Unlike the prefill phase, the decode phase is constrained by memory bandwidth, as it involves GEMV and thin GEMM operations, which inherently limit hardware utilization compared to the more parallelizable GEMMs in the prefill phase.
 Modern inference frameworks such as vLLM (Kwon et al., 2023) and TensorRT-LLM (NVIDIA, 2023) partially remedy this issue by batching multiple decoding requests, improving the computational intensity of the linear layers.
@@
 Furthermore, the computational intensity of the attention operation during decoding is unaffected by batch size, instead requiring a GEMV for each sequence.
+This slowdown becomes more evident at longer sequence lengths, where the attention FLOPs continue to increase in proportion to sequence length while the FLOPs required for linear layers remain constant.
 Grouped query attention (GQA) (Ainslie et al., 2023) converts the operation into a thin GEMM, but it remains memory-bound.
-This slowdown becomes more evident at longer sequence lengths, where the attention FLOPs continue to increase in proportion to sequence length while the FLOPs required for linear layers remain constant.
-5.2 Calculating inference FLOPs
+Multi-head latent attention (MLA) (DeepSeek-AI et al., 2024) further improves the computational intensity during the decode phase, but the trend remains the same.
+5.2. Calculating inference FLOPs
 While previous works (Patel et al., 2024) have evaluated inference performance using time to first token (TTFT) and time per output token (TPOT), these metrics do not facilitate comparisons across different inference stages, model sizes, or sequence lengths.
 To ensure a consistent comparison, we directly compute model FLOPs.
-Following the approach of Ansel et al.
-(2024); Narayanan et al.
-(2021), we calculate only the FLOPs associated with matrix multiplication, as these dominate LLM computations.
-However, unlike previous works, we exclude FLOPs related to autoregressive attention masking, which can be skipped (Dao, 2024).
-This approach aligns more closely with how the KV cache is leveraged when generating new tokens.
+We follow the methodology described in (Ansel et al., 2024; Narayanan et al., 2021), computing only the FLOPs associated with matrix multiplications and excluding those related to autoregressive attention masking, which can often be skipped in practice (Dao, 2024).
+This yields a more faithful representation of the actual compute cost when the KV cache is utilized for incremental decoding.
 Using this method, the FLOPs required for a forward pass of a Llama model with $l$ transformer blocks, hidden size $h$, intermediate size $ah$, head size $d$, head count $H=h/d$, GQA group size $g$, vocabulary size $v$, and input sequence of length $s$ can be calculated as follows:
-$f_{llama}(s)=2sh^{2}l(3a+2+\frac{2}{g})+2s^{2}hl+2vsh$ (2)
-By denoting $A=(3a+2+\frac{2}{g})$ as a constant for each model, the equation can be simplified to the following:
-$f_{llama}(s)=2s(Ah^{2}l+vh)+2s^{2}hl$ (3)
-When the model generates $t$ tokens in a single decoding iteration where $t\ll s$, we can calculate Equation 2 using $s^{\prime}=s+t$ and obtain the approximation:
-$f_{llama}(s+t)-f_{llama}(s)$ (4)
-$\approx 2t(Ah^{2}l+vh)+4sthl$
-From Equation 4, we observe that linear layer computations, including those in the LM head, remain independent of the previous sequence length $s$.
-However, attention computations scale proportionally with $s$.
-In the autoregressive case where $t=1$ and considering a batch of sequences with batch size $b$ and sequence lengths $s_{1},...,s_{b}$, the FLOPs per decoding step can be approximated as follows:
-$2b(Ah^{2}l+vh)+4hl\sum_{i=1}^{b}s_{i}$ (5)
-One challenge in interpreting these results is that only the $2bAh^{2}l$ term, representing linear layers except the LM head, is computed in FP8, whereas the $2bvh$ term for the LM head and the $4hl\sum_{i=1}^{b}s_{i}$ term for attention are computed in BF16.
-Another is that online KV cache dequantization would add non-trivial overhead to Equation 5.
-However, such overheads do not constitute additional model FLOPs as per the definition of MFU in Chowdhery et al.
-(2023).
-A more significant limitation is that each FLOP cannot be executed at full efficiency due to hardware utilization constraints.
-For example, the Gaudi 2 has a peak HBM bandwidth of 2.4 TB/s and a peak FP8 GEMM throughput of 865 TFLOPS, requiring a FLOP/byte ratio (computational intensity, CI) of at least 360 for optimal FP8 execution.
-However, in the decoding phase, a thin GEMM of $(b\times h)\times(h\times ah)$, where $b\ll h$ and $a\geq 1$, results in a CI of approximately $2b$ for FP8 and $b$ for BF16, far below the required intensity for peak throughput on realistic batch sizes $b$.
-Additionally, matrix multiplication units operate with fixed block sizes, meaning that full utilization is only achieved when input dimensions align with hardware-friendly shapes such as multiples of 128 (Lee et al., 2024).
-Furthermore, KV cache computations present a unique bottleneck.
-Increasing batch size does not improve the CI since each sequence in a batch maintains a separate KV cache.
-For a BF16 KV cache using GQA with $g$ groups, the CI is thus limited to $g$ FLOPs/byte.
-Consequently, even with a perfectly optimized kernel, the theoretical throughput of applying the query to the KV cache is capped by the memory bandwidth multiplied by the CI.
-For Llama v3 models with $g=8$ on a Gaudi 2 with a peak 2.4 TB/s HBM, this theoretical ceiling is 19.2 TFLOPS, a small fraction of peak GEMM throughput.
-As attention computations are both inherently memory-bound and scale linearly with sequence length, decoding at longer sequence lengths ultimately converges to the attention throughput, as demonstrated in Figure 1.
-By distinguishing these phases and understanding their computational characteristics, optimization strategies can be tailored to enhance the efficiency of LLM inference, especially during the resource-intensive decode phase.
-5.3 Prefill phase
-Figure 3: Roofline diagram during the prefill phase for batch size 1.
-Throughput improves with longer sequence length until it begins to decline as the proportion of attention computations, which are slower than GEMMs, take up a greater share of the computation.
+(3) $f_{llama}(s)=2sh^{2}l(3a+2+\frac{2}{g})+2s^{2}hl+2vsh.$
+By denoting $A=(3a+2+\frac{2}{g})$ as a model-specific constant, Equation 3 simplifies to:
+(4) $f_{llama}(s)=2s(Ah^{2}l+vh)+2s^{2}hl.$
+When the model generates $t$ tokens in a single decoding iteration where $t\ll s$, we approximate the additional computation by evaluating Equation 3 at $s^{\prime}=s+t$, yielding:
+(5) $f_{llama}(s+t)-f_{llama}(s)$
+$\approx 2t(Ah^{2}l+vh)+4sthl.$
+From Equation 5, we observe that the computational cost of the linear layers, including the LM head, remains independent of the previous sequence length $s$, while the attention-related FLOPs scale proportionally with $s$.
+In the autoregressive setting where $t=1$, and for a batch of $b$ sequences with respective sequence lengths $s_{1},\dots,s_{b}$, the FLOPs required for one decoding step can be approximated as:
+(6) $2b(Ah^{2}l+vh)+4hl\sum_{i=1}^{b}s_{i}.$
+A key observation is that only the first term, $2bAh^{2}l$, corresponding to the linear layers (excluding the LM head), is computed in FP8.
+In contrast, the $2bvh$ term for the LM head and the $4hl\sum_{i=1}^{b}s_{i}$ term for attention are computed in BF16.
+Additionally, online dequantization of the KV cache introduces extra overhead not accounted for in Equation 6.
+However, these overheads do not contribute to the model FLOPs as defined in the MFU metric (Chowdhery et al., 2023).
+A more fundamental limitation is that not every FLOP can be executed at full efficiency due to hardware utilization constraints.
+For instance, the Gaudi 2 features a peak HBM bandwidth of 2.4 TB/s and a peak FP8 GEMM throughput of 865 TFLOPS, which implies that a FLOP/byte ratio (i.e., computational intensity, CI) of at least 360 is required to saturate the compute units.
+However, during the decoding phase, the workload often reduces to a thin GEMM of shape $(b\times h)\times(h\times ah)$, where $b\ll h$ and $a\geq 1$.
+In this case, the resulting CI drops to approximately $2b$ for FP8 and $b$ for BF16, which is far below the required threshold for achieving peak throughput at practical batch sizes.
+Moreover, matrix multiplication units typically operate on fixed-size blocks, and full hardware utilization is only achieved when input dimensions align with hardware-friendly shapes, such as multiples of 128 (Lee et al., 2024).
+Furthermore, KV cache computations introduce a distinct bottleneck.
+Increasing the batch size does not improve CI, as each sequence in the batch maintains an independent KV cache.
+In the case of a BF16 KV cache with GQA using $g$ groups, the CI is bounded by $g$ FLOPs/byte.
+As a result, even with an optimally implemented kernel, the theoretical throughput for applying the query to the KV cache is limited by the product of memory bandwidth and CI.
+For example, in LLaMA v3 models with $g=8$ on a Gaudi 2 with 2.4 TB/s peak HBM bandwidth, the maximum achievable throughput is 19.2 TFLOPS, only a small fraction of the device’s peak GEMM throughput.
+Because attention operations are inherently memory-bound and their FLOPs scale linearly with sequence length, the decoding phase at longer sequence lengths eventually becomes limited by attention throughput.
+Understanding the distinct computational characteristics of each phase allows for more targeted optimization strategies, particularly in improving the efficiency of the decode phase, which dominates end-to-end inference latency.
+5.3. Prefill phase
+We first conducted an analysis of the operational characteristics of the two accelerators in the prefill phase.
+Figure 4 presents a roofline comparison of prefill throughput (TFLOPS) for three models across varying sequence lengths on the H100 and Gaudi 2.
+The results show that the H100 consistently achieves significantly higher throughput than the Gaudi 2 in this phase, reaching up to double the performance of the Gaudi 2 for LLaMA 8B.
+In the prefill phase, attention layers process the entire input sequence at once, enabling large matrix multiplications to be executed in a single step.
+This leads to high utilization of compute engines and results in performance primarily governed by GEMM throughput, rather than memory bandwidth or utilization constraints.
+As model size increases, matrix dimensions scale proportionally with the hidden size, making computations increasingly compute-bound.
+This results in a clear trend of improved prefill throughput for larger models, highlighting the importance of raw compute performance in this phase.
+A similar trend is observed with longer input sequences, where hardware utilization improves until saturation is reached.
+Figure 4.
+Roofline diagram during the prefill phase for batch size 1.
+Throughput improves with longer sequence lengths until they begin to decline as the proportion of attention computations, which are slower than GEMMs, takes up a greater share of the computation.
 We use static FP8 scaling for both the H100 and the Gaudi 2 as optimized dynamic FP8 scaling is unavailable for the Gaudi.
-We first conducted an analysis of the operational characteristics of the two accelerators in the prefill phase.
-Figure 3 compares the prefill TFLOPS achieved for three models across varying sequence lengths, using a roofline diagram for the H100 and Gaudi 2 accelerators.
-The results demonstrate that the H100 achieves significantly higher throughput during the prefill phase than the Gaudi 2.
-For Llama 8B, the H100 achieves double the throughput of the Gaudi 2.
-During the prefill phase, attention layers process long input sequences, enabling large matrix computations to be executed in a single step.
-This maximizes compute engine utilization, leading to high efficiency.
-As a result, performance in this phase is primarily dictated by GEMM throughput, rather than utilization or memory bandwidth constraints.
-As model sizes increase, matrix dimensions grow proportionally with hidden dimensions, making computations more compute-bound.
-This leads to a clear trend of increasing prefill throughput with larger models, underscoring the importance of compute performance in the prefill phase.
-A similar pattern is observed for longer sequence lengths, where utilization improves until reaching saturation.
-Figure 4: Decode throughput comparison between BF16 and FP8 on Llama v3.1 8B Instruct using a batch size of 64 using static FP8 scaling on the Gaudi 2 (left); static and dynamic FP8 scaling on the H100 (right).
+Figure 5.
+Decode throughput comparison between BF16 and FP8 on Llama v3.1 8B Instruct using a batch size of 64 using static FP8 scaling on the Gaudi 2 (left); static and dynamic FP8 scaling on the H100 (right).
 Throughput differences between FP8 and BF16 in the Gaudi 2 are 50% or greater, whereas they are under 25% for the H100.
 Dynamic scaling has higher throughput despite its greater overhead because row-wise GEMM is faster than per-tensor GEMM for small matrices on the H100 (see Table 3).
-5.4 Decode phase
-In Figure 1, we compare decode throughputs between the Gaudi 2 and H100.
-Surprisingly, the Gaudi 2 not only demonstrates superior power efficiency across all sequence lengths but also achieves higher absolute throughput for short sequences below 4K tokens.
-At longer sequence lengths, however, memory bandwidth becomes the primary bottleneck.
-Our measurements show HBM bandwidths of 2.0 TB/s for the Gaudi 2 and 2.6 TB/s for the H100, giving the latter an advantage as the KV cache size increases.
-We also compare decode throughputs for BF16, dynamic FP8 scaling, and static FP8 scaling on the H100 in Figure 4 and find that the throughput gain from FP8 is below 25%.
-In contrast, the gain for the Gaudi 2 approaches 50%.
-To investigate the lack of performance gains in the H100, we analyze its throughput on thin matrices, identifying it as the primary bottleneck.
-Table 4 shows that while FP8 GEMM throughput on the Gaudi 2 is nearly twice that of BF16 GEMM, there is minimal improvement on the H100.
-As shown in Equation 5, the decoding stage at short sequence lengths ($s<h$) is dominated by linear operations, making thin GEMM performance critical.
-Table 4: Thin GEMM throughputs in TFLOPS for the Gaudi 2 and H100 measuring $(M\times K)\times(K\times N)$ GEMMs, where $M$ corresponds to typical batch sizes encountered during inference and $K,N$ represent hidden dimension sizes.
+5.4. Decode phase
+The decode phase has emerged as the primary contributor to overall inference latency in modern LLMs.
+As “reasoning” models (DeepSeek-AI, 2025) often generate thousands of tokens to improve output quality, end-to-end performance is increasingly dominated by the decode phase.
+Consequently, optimizing this phase is critical to reduce inference latency and maximize throughput.
+Figure 5 compares decode throughputs on the H100 across three settings: BF16, dynamic FP8 scaling, and static FP8 scaling.
+On the H100, FP8 yields less than a 25% improvement over BF16, whereas the Gaudi 2 achieves up to a 50% gain.
+These results indicate that the H100 does not fully realize the theoretical FP8 speedup (e.g., double that of BF16) during the decode phase.
+As a result, its FP8 performance is only marginally better than BF16.
+Notably, Gaudi 2 with FP8 achieves comparable decode throughput to the H100, despite having significantly lower peak GEMM throughput.
+5.5. Power capping strategies
+As generative LLM inference tasks become more prevalent, datacenters must contend with substantial power consumption from AI accelerators, leading to high TCO.
+High-performance GPUs can draw significant power, but data centers have limited power infrastructure, making efficient power management crucial.
+One promising strategy is GPU power capping, which imposes a fixed power limit for GPUs.
+This keeps overall power usage predictable and helps prevent out-of-spec conditions, such as rack overheating or hardware damage.
+To observe the performance of each GPU in constrained conditions, we measured the throughput of the Gaudi 2 and H100 in a constrained power environment.
+Given that much existing infrastructure in production datacenters has been built around the A100, with infrastructure commonly optimized for its 400W TDP, we perform our evaluations under the same power budget to ensure a fair and compatible comparison.
+As noted in Patel et al.
+(2024), the decode phase does not show throughput deterioration with power capping at 400W.
+Figure 3 compares the theoretical roofline with the measured TFLOPS under a 400W power constraint.
+We show only the 400W measurements as the measurements without power capping are nearly indistinguishable during decoding.
+Note that while the H100 exhibits a much higher theoretical throughput, the Gaudi 2 achieves higher measured performance in several scenarios.
+However, both devices currently support per-GPU power capping only, rather than per-rack power capping.
+Per-rack power capping offers a compelling solution by preserving the benefits of power management while allowing each GPU greater flexibility in power consumption.
+When power is constrained at the individual GPU level, no device can exceed its fixed limit, potentially leaving unused power headroom untapped.
+In contrast, per-rack capping treats the total power budget of each rack as a collective resource, letting one GPU utilize surplus capacity if another GPU draws less than its share.
+This approach can lead to more efficient resource usage and higher overall performance.
+5.6. Thin GEMM with FP8 quantization
+Figure 6.
+Thin GEMM MFU comparison between the Gaudi 2 and H100 for both BF16 and FP8.
+The Gaudi 2 maintains a similar MFU for BF16 and FP8 of similar shapes while there is a noticeable drop for the H100.
+The MFU differences at the same shapes are enough to provide superior TFLOPS for the Gaudi 2 over the H100 as shown in Table 6.
+Thin GEMM is a critical workload that significantly impacts overall inference latency.
+Although LLM sizes continue to scale, their hidden dimensions grow at a much slower rate.
+For example, Llama 8B has a hidden size of 4096, whereas Llama v3.3 70B increases it only to 8192.
+As shown in Equation 6, the decoding stage at short sequence lengths ($s<h$) is dominated by linear operations.
+The growing adoption of fine-grained Mixture-of-Experts (MoE) models (DeepSeek-AI, 2024; Meta, 2025) further contributes to this trend, as a larger number of experts reduces the average number of activations assigned to each expert during batched decoding.
+Additionally, tensor parallelism (TP), which is commonly employed to reduce latency and memory overhead, further partitions model weight matrices into smaller submatrices.
+As a result, the efficiency of small matrix operations has become a key factor in optimizing LLM inference performance.
+Unlike large matrix-matrix multiplications, thin GEMMs fail to fully utilize compute units, making it difficult to achieve peak performance.
+Moreover, the inefficiency that derives from hardware utilization increases significantly as low data precision is adopted.
+Theoretically, FP8 offers up to double the throughput of BF16 or FP16.
+However, even if the compute units themselves are capable of fast execution, suboptimal input/output pipelining or memory access patterns can severely bottleneck overall throughput.
+Therefore, to operate datacenters efficiently, it is crucial to understand how specific hardware platforms perform in scenarios involving thin GEMM and low-precision computation.
+Based on the insight about the thin GEMM utilization difficulties, we examine the behavior of the Gaudi 2 and H100 under a combination of thin GEMM workloads and FP8 precisions.
+Table 6 presents the GEMM TFLOPS in a setting where the dimension of the matrix $M$ is significantly smaller than $K$ and $N$, which expresses the thin GEMM condition.
+Compared to its performance on large GEMMs, the H100 shows notably lower throughput than its specifications under this thin GEMM scenario.
+FP8 and BF16 throughputs for the H100 are nearly identical, failing to reflect the expected performance gains from lower precision.
+In contrast, the Gaudi 2 exhibits an FP8 throughput approximately double that of BF16, which aligns closely with theoretical expectations.
+Table 6.
+Thin GEMM throughputs (in TFLOPS) for the Gaudi 2 and H100 measuring $(M\times K)\times(K\times N)$ GEMMs, where $M$ corresponds to typical batch sizes during inference and $K,N$ represent hidden dimension sizes.
 Row-wise scaling is used for FP8 GEMMs.
 Throughput scales linearly with $M$ on both devices, but the Gaudi 2 consistently outperforms the H100, even for BF16.
-These results highlight the superior efficiency of the Gaudi 2 for thin GEMMs, crucial for LLM decoding.
+Results highlight the superior efficiency of the Gaudi 2 for thin GEMMs, crucial for LLM decoding.
 GEMM TFLOPS Gaudi 2 H100
@@
 (64, 4096, 4096) 144.5 253.4 133.3 133.9
-Figure 5: Thin GEMM MFU comparison between the Gaudi 2 and H100 for both BF16 and FP8.
-The Gaudi 2 maintains a similar MFU for BF16 and FP8 of similar shapes while there is a noticeable drop for the H100.
-The MFU differences at the same shapes are enough to provide superior TFLOPS for the Gaudi 2 over the H100 as shown in Table 4.
-6 Conclusion
-In this study, we provide a comprehensive analysis of FP8, an emerging numerical format that remains underexplored.
-We examine various FP8 scaling factor configurations and analyze LLM inference workloads based on the constraints of real-world data center environments.
-We make a clear exposition of the key constraints in LLM inference, also providing a framework to make precise calculations for inference throughput.
-Our analysis of the two phases of LLM inference highlights the strong dependence of LLM inference on hardware characteristics.
-However, beyond hardware development, software optimizations that fully leverage accelerator capabilities are equally essential to maximizing performance.
-One key limitation of the Gaudi is its limited software support and integration.
-For instance, many kernel fusion techniques for dynamic FP8 scaling are currently unavailable on the Gaudi.
-Additionally, the decode throughputs reported in our study relied on static memory allocation as the paged attention implementations remained uncompetitive despite recent optimizations (Lee et al., 2024).
-Even for static memory allocation, the observed attention throughputs are lower than what would be expected from the hardware specifications.
-Furthermore, while not a primary focus of this work, the lack of KV cache quantization support present challenges for deployment.
-Despite these concerns, the growing adoption of inference-time scaling via chain-of-thought (CoT) (Wei et al., 2022) prompting suggests that LLM workloads will increasingly be dominated by decoding.
-Emerging techniques that decouple the prefill and decode phases (Patel et al., 2024; Zhong et al., 2024) enable hardware to be allocated to the tasks for which they are best suited.
-Given our findings that the Gaudi 2 demonstrates superior decoding efficiency for FP8, this has potential implications for cloud service providers seeking cost-effective inference solutions.
-Appendix A Appendix
-Table 5: MMLU result for per-row, per-tensor options
-  model Rowwise Scaling accuracy
-Activation Weight
-Llama-3.2-1B 35.12%
-34.69%
-34.36%
-34.65%
-Llama-3.2-3B 53.90%
-53.92%
-53.83%
-53.75%
-Llama-3.1-8B 62.95%
-62.92%
-62.80%
-62.90%
-Llama-3.1-70B 75.25%
-75.23%
-75.32%
-75.39%
-Table 6: MMLU 5-shot CoT for stochastic rounding using E4M3.
-  Model Stochastic Rounding MMLU accuracy
-Activation Weight
-Llama-3.2-1B 33.94%
-34.31%
-35.20%
-34.65%
-Llama-3.2-3B 54.09%
-53.95%
-54.41%
-53.75%
-Llama-3.1-8B 62.24%
-62.43%
-62.55%
-62.90%
-Llama-3.1-70B 75.05%
-75.18%
-75.30%
-75.39%
-Table 7: Throughput and Power Measurement of Llama Instruct Models
-  Model Name Mode Batch Size Seq Len Power(W) Prefill TFLOPS/HPU
-Llama-3.2-1B-Instruct FP8-Static 4 512 509 412
-Llama-3.2-1B-Instruct FP8-Static 4 1024 509 411
-Llama-3.2-3B-Instruct FP8-Static 4 512 498 518
-Llama-3.2-3B-Instruct FP8-Static 4 1024 511 537
-Llama-3.1-8B-Instruct FP8-Static 4 512 471 588
-Llama-3.1-8B-Instruct FP8-Static 4 1024 436 595
-Table 8: Gaudi 2 throughput in TFLOPS for scaled FP8 GEMM for square matrices of the given size, excluding quantization overhead.
-Measured throughput relative to the peak FP8 throughput (865 TFLOPS) is included in parentheses.
-Hardware acceleration is only available for per-tensor scaling.
-Type Size Per-Row Per-Tensor HW Accel.
-E4M3 1K 494 (57.1%) 494 (57.1%) 494 (57.1%)
-2K 506 (58.5%) 641 (74.1%) 641 (74.2%)
-4K 735 (84.9%) 796 (92.1%) 801 (92.6%)
-8K 742 (85.7%) 822 (95.0%) 852 (98.4%)
-E5M2 1K 306 (35.4%) 494 (57.1%) 493 (57.0%)
-2K 506 (58.5%) 642 (74.2%) 642 (74.2%)
-4K 735 (84.9%) 802 (92.7%) 802 (92.7%)
-8K 726 (83.9%) 825 (95.4%) 825 (95.4%)
-Table 9: H100 throughput in TFLOPS for scaled FP8 GEMM for square matrices of the given size, excluding quantization overhead.
-Measured throughput as a ratio of peak FP8 throughput (1989.9 TFLOPS) is included in parentheses.
-We only include results for E4M3 because we were unable to find libraries exposing low-level E5M2 GEMM APIs on NVIDIA GPUs.
-Accum. Size Per-Row Per-Tensor
-FP32 1K 217 (11.0%) 186   (9.4%)
-2K 299 (15.1%) 840 (42.4%)
-4K 362 (18.3%) 1099 (55.5%)
-8K 396 (20.0%) 1300 (65.7%)
-Fast 1K 237 (12.0%) 147   (7.4%)
-2K 810 (40.9%) 896 (45.3%)
-4K 1136 (57.4%) 1205 (60.9%)
-8K 1123 (56.8%) 1388 (70.1%)
+These differences in FP8 thin GEMM performance are likely attributed to the architectural differences of the hardware platforms.
+In the case of NVIDIA Hopper, details of the Tensor Core architecture remain undisclosed, making it difficult to analyze the actual implementation.
+We expect that future disclosure of internal implementation details would enable a more systematic investigation.
+In contrast, Gaudi2 provides public documentation that contains their Matrix Multiplication Engine (MME) architecture, which is an output-stationary systolic array.
+Figure 7 depicts the differences between the MME architectures of the Gaudi2 and those of other commercial hardware solutions.
+Unlike other commercial hardware offerings that employ multiple small compute units, Gaudi2 integrates a single large MME.
+In architectures that rely on several small units, each unit consumes input data in every cycle, incurring significant overhead in unit-to-unit data movement.
+In contrast, one large MME consumes a much smaller amount of input data per cycle and requires less expensive ALU-to-ALU communication within a whole MME.
+The effect of communication cost leads to notable MFU differences, as shown in Figure 6.
+The data movement latency is effectively hidden in compute-bound conditions, but it becomes critical in memory-bound thin GEMM operations, where communication overhead can dominate performance and markedly degrade overall performance.
+Furthermore, Figure 8 illustrates the reconfigurable MME sizes of Gaudi 2.
+While the Gaudi2 architecture contains two MMEs with size $256\times 256$, they can be folded and integrated in different sizes with a smallest width 128.
+The size of the MME is optimized at run time via the Gaudi Graph Compiler.
+This flexibility allows the hardware to adapt to a wide range of GEMM dimensions, offering substantial potential for improving LLM inference efficiency.
+5.7. Exponential function optimization
+While the Gaudi 2 outperforms the H100 in the LLM decode phase for short sequence lengths, its throughput gains diminish at longer sequence lengths.
+We find that this decline is primarily due to the absence of the dedicated Special Function Units (SFUs, also referred to as multi-function units, or MUFUs) in the Gaudi 2 & 3 to accelerate the exponential calculations in the softmax.
+During typical LLM inference, SFUs perform crucial tasks.
+Although GEMM operations that leverage tensor cores dominate LLM workloads, specialized functions such as softmax and SiLU are also necessary.
+For decoding workloads, the ratio of exponential computations relative to GEMM computations increase in proportion to the sequence length.
+As a result, the softmax operation becomes a bottleneck for decoding at long sequence lengths.
+GPUs equipped with SFUs can process exponential functions on SFUs while simultaneously executing GEMMs on tensor cores, effectively parallelizing the workload and boosting overall performance.
+In contrast, both the Gaudi 2 & 3 have a disadvantage in that they lack SFUs and must handle exponential computations via their TPC cores.
+They also have a lower ratio of vector core compute relative to matrix computation units compared to the H100.
+Peak TPC core BF16 throughput is 11 TFLOPS for the Gaudi 2 and 28.7 TFLOPS for the Gaudi 3, whereas the H100 has 133.8 TFLOPS BF16 CUDA core throughput and includes separate SFU cores.
+This implies that despite the notable efficiency of thin GEMMs in the Gaudi 2, exponential function calculations emerge as the main bottleneck for long sequences, as the softmax computation scales $O(BS)$ for batch size $B$ and sequence length $S$ during decoding.
+The Gaudi 2 incorporates an embedded look-up table (LUT), which could theoretically store precomputed special-function values via piecewise linear approximation.
+If the LUT offers sufficiently low access latency and can hold a large enough range of values, it may be a valid substitute for SFUs.
+However, Gaudi 2 currently lacks software support for programming the LUT.
+Although it has strong potential to improve performance with long sequences, the limited software stack constrains its practical effectiveness.
+Figure 7. Gaudi2 employs one large systolic array unit for each MME, unlike the many other commercial hardware employs many small units.
+Figure 8. Dynamically reconfigurable MME architecture on Gaudi for target GEMM optimization.
+6. Throughput shifts the TCO balance
+As described in Section 2, the TCO varies significantly depending on factors such as data precision, workload characteristics, and batch size.
+Figure 9 illustrates how the TCO ratio changes under varying conditions, particularly focusing on two key parameters: the throughput ratio and the server cost ratio.
+In our TCO model, system A represents the Gaudi 2 and system B represents the H100.
+In Figure 9, movement along the x-axis reflects the server cost ratio, which is determined by factors such as device procurement cost, infrastructure constraints, and other datacenter-specific considerations.
+This ratio varies across CSPs depending on their individual deployment environments.
+Therefore, the dominant factor influencing the TCO ratio is the throughput ratio ($R_{Th}$).
+In the LLM decoding phase, which dominates overall inference latency, the Gaudi 2 achieves nearly 150% the throughput of BF16 when using FP8 quantization, while the H100 shows little to no performance gain.
+As a result, $R_{Th}$ increases, and the TCO ratio shifts upward along the y-axis, approaching the green region in the figure, indicating higher cost efficiency for the Gaudi 2.
+Conversely, in scenarios involving long sequences, the Gaudi 2 suffers from the softmax bottleneck due to its lack of a dedicated SFU or an alternative workaround.
+This lowers the throughput ratio, shifting the TCO ratio downward along the y-axis toward the red region, which implies reduced cost efficiency for the Gaudi 2.
+Ultimately, the actual position on the TCO map depends on the specific conditions of each datacenter and on the specific operations required by the CSP.
+Therefore, accurate evaluation requires a detailed analysis of the operational environment and workload profile unique to each CSP.
+Figure 9. Visualization of TCO Variation by Server Cost and Throughput Ratios.
+7. Conclusion
+In this study, we present a scalable framework for comparing AI accelerator options in terms of datacenter TCO.
+Using our model, we examine two key factors that influence throughput and are directly connected to the TCO: numerical precision and LLM inference workload characteristics.
+We evaluate two representative next-generation accelerators, the NVIDIA H100 and Intel Gaudi 2, and analyze their behavior under FP8 quantization.
+However, our study has several limitations.
+First, our analysis is based on a case study focused specifically on Llama-based models in the context of inference.
+As such, the results may not directly generalize to other workloads, such as training.
+Second, we adopt FP8 precision to reflect anticipated trends in future-generation models.
+However, the applicability of our findings depends on the extent of hardware support for FP8, which remains a limitation.
+While our study provides valuable insights into cost-performance trade-offs, additional research is required to generalize the results to full-scale datacenter environments with diverse infrastructure, workload mixes, and deployment strategies.
+We hope that the insights provided in this study help inform deployment decisions and guide future accelerator design efforts aimed at improving the TCO of LLM inference in datacenter environments.